### v3

# Density of irreducible operators in the trace-class norm

Abstract.
In 1968, Paul Halmos initiated the research on density of the set of irreducible operators on a separable Hilbert space. Through the research, a long-standing unsolved problem inquires: is the set of irreducible operators dense in $\mathcal{B}(\mathcal{H})$ with respect to the trace-class norm topology? Precisely, for each operator $T$ in $\mathcal{B}(\mathcal{H})$ and every $\varepsilon>0$, is there a trace-class operator $K$ such that $T+K$ is irreducible and $\|K\|_{1}<\varepsilon$?
For $p>1$, to prove the $\|\cdot\|_{p}$-norm density of irreducible operators in $\mathcal{B}(\mathcal{H})$, a type of Weyl-von Neumann theorem effects as a key technique. But the traditional method fails for the case $p=1$, where by $\|\cdot\|_{p}$-norm we denote the Schatten $p$-norm.
In the current paper, for a large family of operators in $\mathcal{B}(\mathcal{H})$, we give the above long-term problem an affirmative answer. The result is derived from a combination of techniques in both operator theory and operator algebras. Moreover, we discover that there is a strong connection between the problem and another related operator-theoretical problem related to type $\mathrm{II}_{1}$ von Neumann algebras.
Key words and phrases:
Schatten $p$-norm, von Neumann algebra, irreducible operator, trace-class operator.
2020 Mathematics Subject Classification:
Primary 47C15
This article was partly supported by Tianyuan Mathematics Research Center. Chunlan Jiang and Junsheng Fang are partly supported by the Hebei Natural Science Foundation (No.A2023205045)
Rui Shi, Minhua Ma, and Tianze Wang are partly supported by NSFC (No.12271074), the Fundamental Research Funds for the Central Universities (No.DUT23LAB305), and a research fund from Math Department of HEBNU for visiting scholars.
1. Introduction
We present the long-standing problem as follows.
Problem A.
Is the set of irreducible operators dense in $\mathcal{B}(\mathcal{H})$ in the trace-class norm?
Throughout this paper, let $\mathcal{H}$ be a complex separable infinite-dimensional Hilbert space and $\mathcal{B}(\mathcal{H})$ the algebra of all bounded linear operators on $\mathcal{H}$. Recall that an operator $T$ in $\mathcal{B}(\mathcal{H})$ is irreducible if $T$ has no nontrivial reducing subspaces. That is to say, if $P$ is a projection (i.e., $P=P^{*}=P^{2}$) in $\mathcal{B}(\mathcal{H})$ such that $PT=TP$, then $P=0$ or $P=I$. By definition, the irreducibility of operators is invariant up to unitary equivalence.
Problem A can be traced back to a result of Paul Halmos in [9]. In the following, we briefly recall Halmos’ original result about irreducible operators, contributions of other people related to Problem A, main techniques applied before, and the reason that the Weyl-von Neumann theorem fails to effect towards the solution of Problem A.
1.1. Density problem of irreducible operators in $\mathcal{B}(\mathcal{H})$
By definition, irreducible operators can be viewed as atoms to construct operators in $\mathcal{B}(\mathcal{H})$. In this sense, it is natural to explore how large the set of irreducible operators is. In the operator norm topology, Paul Halmos proved that irreducible operators form a dense $G_{\delta}$ subset of $\mathcal{B}(\mathcal{H})$ in [9]. Later, Heydar Radjavi and Peter Rosenthal gave a short proof in [21]. It turns out that, on considering the operator norm density of the set of irreducible operators, one needs the classical form of the spectral theorem for bounded self-adjoint operators and matrix-construction techniques.
In the last paragraph of [9, Section 1], Ronald Douglas observed that by virtue of the Weyl-von Neumann theorem, Halmos’ density theorem is also true in the sense of Hilbert-Schmidt approximation. To improve the above result with the Schatten $p$-norm, one needs a type of the Weyl-von Neumann theorem for self-adjoint operators as a key technique. In the following part, we denote by $\|\cdot\|_{p}$-norm the Schatten $p$-norm for $p\geq 1$. Note that the Schatten $2$-norm is the Hilbert-Schmidt norm, while the Schatten $1$-norm is the trace-class norm.
The classical Weyl-von Neumann theorem for self-adjoint operators due to Hermann Weyl [31] and John von Neumann [18] states that every self-adjoint operator in $\mathcal{B}(\mathcal{H})$ is diagonalizable up to an arbitrarily small Hilbert-Schmidt perturbation.
In [16], Shige Toshi Koruda improved the classical Weyl-von Neumann theorem by proving that every self-adjoint operator in $\mathcal{B}(\mathcal{H})$ is diagonalizable up to an arbitrarily small $\Phi$-norm perturbation, where by $\Phi$-norm we denote a unitarily invariant norm not equivalent to the trace-class norm. Note that, for every $p>1$, the $\|\cdot\|_{p}$-norm serves as a candidate for such a unitarily invariant norm. According to the Weyl-von Neumann-Kuroda theorem in [16] and techniques of H. Radjavi and P. Rosenthal developed in [21], Domingo Herrero proved in [10, Lemma 4.33] that the set of irreducible operators is $\|\cdot\|_{p}$-norm dense in $\mathcal{B}(\mathcal{H})$ for every $p>1$.
1.2. Schatten 1-norm perturbations of self-adjoint operators
It seems that the research on the density of the set of irreducible operators with respect to the $\|\cdot\|_{p}$-norm would be intact if the Weyl-von Neumann theorem holds for $p=1$. But, with respect to the trace-class norm, a large family of self-adjoint operators fails to be diagonalizable up to trace-class perturbations. According to [15, 20], Tosio Kato and Marvin Rosenblum (independently) showed that, up to unitary equivalence, the ((((spectrally)))) absolutely continuous part of a self-adjoint operator in $\mathcal{B}(\mathcal{H})$ is stable under a self-adjoint trace-class perturbation. It is also worth mentioning that, in [2], Richard Carey and Joel Pincus showed that a purely singular self-adjoint operator in $\mathcal{B}(\mathcal{H})$ is a small trace-class perturbation of a diagonal operator.
By the Kato-Rosenblum theorem, the method in Herrero’s proof of [10, Lemma 4.33] with the Weyl-von Neumann-Kuroda theorem fails to effect for $\|\cdot\|_{1}$-norm. Therefore, to investigate Problem A, one has to develop new methods and techniques.
1.3. Main theorem and an outline of the proof
In this paper, for a large family of operators in $\mathcal{B}(\mathcal{H})$, we answer Problem A affirmatively. Based on Proposition 5.2.7 and the discussion in Section 5, we propose a conjecture related to single generators of type $\mathrm{II}_{1}$ von Neumann algebras.
Conjecture 1.3.1.
Assume that $A$ and $B$ are self-adjoint operators generating a type $\mathrm{II}_{1}$ von Neumann algebra $\mathcal{M}$ acting on a complex separable Hilbert space $\mathcal{H}$.
Then for every $\varepsilon>0$, there exists a trace-class operator $K$ in $\mathcal{B}(\mathcal{H})$ with $\|K\|_{1}\leq\varepsilon$ such that $(A+iB)+K$ is a direct sum of at most countably many irreducible operators.
With 1.3.1, we prove the following result in this paper.
Theorem 1.3.2 (Main Theorem).
Suppose that Conjecture 1.3.1 is true. Then the set of irreducible operators is dense in $\mathcal{B}(\mathcal{H})$ with respect to the trace-class norm topology.
Before introducing the method to prove the Main Theorem, we briefly recall the type decomposition theorem for von Neumann algebras. For a von Neumann algebra $\mathcal{M}$ acting on $\mathcal{H}$ (separable and infinite-dimensional), by [14, Theorem 6.5.2], there exist central projections $P_{{\rm I}_{n}}$ ($n\geq 1$), $P_{{\rm I}_{\infty}}$, $P_{{\rm II}_{1}}$, $P_{{\rm II}_{\infty}}$, and $P_{{\rm III}}$, with sum $I$, such that $\mathcal{M}$ can be expressed as a (possibly finite) direct sum of von Neumann algebras in the form
$\mathcal{M}=\left(\bigoplus^{\infty}_{n=1}\mathcal{M}P_{{\rm I}_{n}}\right)\oplus\mathcal{M}P_{{\rm I}_{\infty}}\oplus\mathcal{M}P_{{\rm II}_{1}}\oplus\mathcal{M}P_{{\rm II}_{\infty}}\oplus\mathcal{M}P_{{\rm III}},$ (1.1)
where $\mathcal{M}P_{{\rm I}_{n}}$ is of type ${\rm I}_{n}$ or $P_{{\rm I}_{n}}=0$, $\mathcal{M}P_{{\rm I}_{\infty}}$ is of type ${\rm I}_{\infty}$ or $P_{{\rm I}_{\infty}}=0$, $\mathcal{M}P_{{\rm II}_{1}}$ is of type ${\rm II}_{1}$ or $P_{{\rm II}_{1}}=0$, $\mathcal{M}P_{{\rm II}_{\infty}}$ is of type ${\rm II}_{\infty}$ or $P_{{\rm II}_{\infty}}=0$, and $\mathcal{M}P_{{\rm III}}$ is of type ${\rm III}$ or $P_{{\rm III}}=0$. The reader is referred to [14, Definition 6.5.1] for different types of von Neumann algebras. For the sake of simplicity, we denote by $\mathcal{M}_{{\rm I}_{f}}$ the direct sum $\bigoplus^{\infty}_{n=1}\mathcal{M}P_{{\rm I}_{n}}$, which is sometimes referred to as a finite type I von Neumann algebra. Also, denote by $\mathcal{M}_{\infty}$ the direct sum $\mathcal{M}P_{{\rm I}_{\infty}}\oplus\mathcal{M}P_{{\rm II}_{\infty}}\oplus\mathcal{M}P_{{\rm III}}$, which is a properly infinite von Neumann algebra (see [14, Definition 6.3.1]). Thus, we can rewrite the decomposition in (1.1) as
$\mathcal{M}=\mathcal{M}_{{\rm I}_{f}}\oplus\mathcal{M}P_{{\rm II}_{1}}\oplus\mathcal{M}_{\infty}.$ (1.2)
We briefly introduce the steps to prove the Main Theorem as follows.
Step 1. For an operator $T$ in $\mathcal{B}(\mathcal{H})$, we write $T=A+iB$, where $A$ and $B$ are self-adjoint operators. By Lemma 4.1.2, there exists an arbitrarily small self-adjoint trace-class operator $K_{A}$ such that $A+K_{A}$ and $B$ are in the form
$A+K_{A}:=\left(\begin{array}[]{c|ccc}\alpha&0&0&0\\ \hline\cr 0&A_{1}&0&0\\ 0&0&A_{2}&0\\ 0&0&0&A_{\infty}\\ \end{array}\right)\begin{array}[]{l}\operatorname{ran}E\\ \mathcal{H}_{1}\\ \mathcal{H}_{2}\\ \mathcal{H}_{\infty}\\ \end{array},\quad B:=\left(\begin{array}[]{c|ccc}\beta&\xi^{*}_{1}&\xi^{*}_{2}&\xi^{*}_{\infty}\\ \hline\cr\xi_{1}&B_{1}&0&0\\ \xi_{2}&0&B_{2}&0\\ \xi_{\infty}&0&0&B_{\infty}\\ \end{array}\right)\begin{array}[]{l}\operatorname{ran}E\\ \mathcal{H}_{1}\\ \mathcal{H}_{2}\\ \mathcal{H}_{\infty}\\ \end{array}.$ (1.3)
The notation in (1.3) are explained as follows.
  1. (1)
    $\alpha$ is an isolated eigenvalue of $A+K_{A}$ with multiplicity $1$ and $\beta\in\mathbb{R}$.
  2. (2)
    $E$ is the rank-one spectral projection for $A+K_{A}$ corresponding to $\{\alpha\}$.
  3. (3)
    $\xi_{j}$ is a vector in a column form and $\xi^{*}_{j}$ is the conjugate vector of $\xi_{j}$ in a row form for $j=1,2,\infty$.
  4. (4)
    Write $X:=(I-E)(T+K_{A})(I-E)$ to be an operator on $\text{ran}(I-E)$. According to the decomposition mentioned in (1.2), in the von Neumann algebra $W^{*}(X)$, there exist (mutually orthogonal) central projections $E_{1}$, $E_{2}$, and $E_{\infty}$, with sum $I-E$, such that $W^{*}(X)$ can be expressed as
    $W^{*}(X)=W^{*}(X_{1})\oplus W^{*}(X_{2})\oplus W^{*}(X_{\infty}),$ (1.4)
    where $X_{j}:=XE_{j}$ for $j=1,2,\infty$, and $W^{*}(X_{1})$ is of finite type I or $E_{1}=0$, $W^{*}(X_{2})$ is of type ${\rm II}_{1}$ or $E_{2}=0$, and $W^{*}(X_{\infty})$ is properly infinite or $E_{\infty}=0$. Correspondingly, $W^{*}(X_{j})$ acts on $\mathcal{H}_{j}:=\operatorname{ran}E_{j}$ for $j=1,2,\infty$.
  5. (5)
    Write $A_{j}:=\operatorname{Re}X_{j}$ and $B_{j}:=\operatorname{Im}X_{j}$ for $j=1,2,\infty$.
Step 2. In (1.3), if $E_{1}\neq 0$, then we prove in Proposition 3.2.7 that there is an arbitrarily small trace-class operator $K_{1}$ in $\mathcal{B}(\mathcal{H}_{1})$ such that $(A_{1}+iB_{1})+K_{1}$ is a direct sum of at most countably many irreducible operators. The whole Section 3 is prepared for Proposition 3.2.7.
Step 3. For every properly infinite von Neumann algebra, we prove in Lemma 4.2.4 that the set of generating vectors is dense. In Proposition 4.2.6 we develop a method to construct irreducible operators, which serves for the proof of Theorem 1.3.2. If $E_{1}=0$, $E_{2}=0$, and $E_{\infty}\neq 0$, then by Proposition 4.2.6, we obtain Proposition 4.2.7, which is a special case of Theorem 1.3.2.
Step 4. Assume that 1.3.1 is true. Based on the above steps, we prove that $A+iB$ can be expressed as an irreducible operator on $\mathcal{H}$ up to an arbitrarily small trace-class perturbation.
Above all, to prove the Main Theorem, we need to apply operator approximation theory with respect to the trace-class norm, single generator techniques in $\mathcal{B}(\mathcal{H})$, and techniques of von Neumann algebras.
The paper is organized as follows. In Section 2, we prepare some useful tools to construct irreducible operators. Also, we recall some basic knowledge about the theory of direct integrals. In Section 3, we prove in Proposition 3.2.7 that each operator in a finite type I von Neumann algebra can be written as a direct sum of at most countably many irreducible operators up to an arbitrarily small trace-class perturbation. In Section 4, we develop perturbation techniques to construct irreducible operators with respect to the form as in (4.6) and single generators of properly infinite von Neumann algebras. As an application of Proposition 4.2.6, we prove in Proposition 4.2.7 that each operator as in (4.4) can be written as an irreducible operator up to an arbitrarily small trace-class perturbation, which is a special case of Theorem 1.3.2. In Section 5, we start by developing perturbation techniques for single generators of type $\mathrm{II}_{1}$ von Neumann algebras in Proposition 5.1.2. Then in Subsection 5.2 we remove the restriction in Proposition 5.1.2 on the underlying Hilbert space for singly generated type ${\rm II}_{1}$ factors. For self-adjoint operators $A$ and $B$ generating a type ${\rm II}_{1}$ factor with $W^{*}(A)$ non-diffuse, we prove in Proposition 5.2.2 that there is an arbitrarily small trace-class operator $K$ such that $(A+iB)+K$ is irreducible on $\mathcal{H}$. On the other hand, if $W^{*}(A)$ is diffuse, then, in addition, we assume that $W^{*}(A+iB)$ is a hyperfinite type ${\rm II}_{1}$ factor with $W^{*}(A)$ a Cartan subalgebra. With the setting, we prove in Proposition 5.2.7 that there is an arbitrarily small self-adjoint trace-class operator $K$ on $\mathcal{H}$ such that $A+i(B+K)$ is a direct sum of irreducible operators. In Example 5.2.8, we observe a similar result for $W^{*}(A+iB)$ being a class of interpolated free group factors. Thus, by the results we obtained in Section 5, it is natural to propose 1.3.1. Finally, in Section 6, we finish the proof of Theorem 1.3.2.
2. Preliminaries
For a complex separable Hilbert space $\mathcal{H}$, we write $\mathcal{H}^{(n)}$ for the orthogonal direct sum of $n$ copies of $\mathcal{H}$. We also write $\mathcal{H}\oplus\mathcal{H}$ for $\mathcal{H}^{(2)}$. For every operator $T$ in $\mathcal{B}(\mathcal{H})$ and every vector $\xi_{1}\oplus\cdots\oplus\xi_{n}$ in $\mathcal{H}^{(n)}$, denote by $T^{(n)}$ the operator
$T^{(n)}(\xi_{1}\oplus\cdots\oplus\xi_{n}):=T\xi_{1}\oplus\cdots\oplus T\xi_{n}.$
By $W^{*}(T)$ we denote the von Neumann algebra generated by $T$ acting on $\mathcal{H}$. For a subset $\mathcal{S}$ of $\mathcal{B}(\mathcal{H})$, write $\mathcal{S}^{\prime}:=\{X\in\mathcal{B}(\mathcal{H})\ |\ XS=SX,\text{ for every }S\in\mathcal{S}\}$ to be the commutant of $\mathcal{S}$ in $\mathcal{B}(\mathcal{H})$.
2.1. Classical tools to construct irreducible operators
With irreducible operators, spectral theory and matrix techniques, we can construct more irreducible operators.
Lemma 2.1.1.
Let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be complex separable Hilbert spaces. Suppose that $A$ and $B$ are self-adjoint operators in $\mathcal{B}(\mathcal{H}_{1}\oplus\mathcal{H}_{2})$ of the form
$A:=\begin{pmatrix}A_{1}&0\\ 0&A_{2}\end{pmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\end{array}\quad\text{ and }\quad B:=\begin{pmatrix}B_{11}&B_{12}\\ B_{21}&B_{22}\end{pmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\end{array}$
satisfying
  1. $(1)$
    $A_{j}+iB_{jj}$ is irreducible on $\mathcal{H}_{j}$ for $j=1,2$;
  2. $(2)$
    $\sigma\left(A_{1}\right)\cap\sigma\left(A_{2}\right)=\varnothing$;
  3. $(3)$
    $B_{12}\neq 0$.
Then the operator $A+iB$ is irreducible in $\mathcal{B}(\mathcal{H}_{1}\oplus\mathcal{H}_{2})$.
Proof.
Let $Q$ be a projection in $\mathcal{B}(\mathcal{H}_{1}\oplus\mathcal{H}_{2})$ commuting with $A+iB$. Employing the function calculus for $A$ and the hypothesis $\sigma\left(A_{1}\right)\cap\sigma\left(A_{2}\right)=\varnothing$, we have that the projection $P_{1}$ onto $\mathcal{H}_{1}$ lies in the unital $C^{*}$-algebra $C^{*}(A)$. Thus, $P_{1}\in W^{*}(A)$. It follows that $P_{1}Q=QP_{1}$. This means that the projection $Q$ can be expressed in the form
$Q=\begin{pmatrix}Q_{1}&0\\ 0&Q_{2}\end{pmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\end{array}.$
It follows that $Q_{j}$ commutes with $A_{j}+iB_{jj}$ for $j=1,2$. This combining with the irreducibility of each $A_{j}+iB_{jj}$ yields that each $Q_{j}$ is trivial on $\mathcal{H}_{j}$. Since $B_{12}\neq 0$ and $Q_{1}B_{12}=B_{12}Q_{2}$, we have that the projection $Q_{1}+Q_{2}$ is trivial. This completes the proof. ∎
To avoid confusion in later sections, for two vectors $e$ and $f$ in $\mathcal{H}$, we denote by $e\otimes f$ a tensor product vector in $\mathcal{H}\otimes\mathcal{H}$ and by $e\hat{\otimes}f$ we denote a rank-one operator acting on $\mathcal{H}$ defined in the form
$(e\hat{\otimes}f)(h):=\langle h,f\rangle e,\quad\forall\ h\in\mathcal{H}.$ (2.1)
When no confusion can arise, for a vector $e$ in $\mathcal{H}$, we denote by $\|e\|:=\langle e,e\rangle^{\frac{1}{2}}$ the vector norm of $e$ and denote by $\operatorname{Tr}$ the standard trace on the set of trace-class operators. In particular, if $e$ is a unit vector, then the rank-one operator $e\hat{\otimes}e$ is a projection and $\operatorname{Tr}(e\hat{\otimes}e)=1$.
We define $\|\cdot\|_{1}$-diagonalizable self-adjoint operators as follows. Suppose that $A$ is a self-adjoint operator in $\mathcal{B}(\mathcal{H})$. If for every $\varepsilon>0$ there is a self-adjoint trace-class operator $K$ satisfying $\|K\|_{1}<\varepsilon$ such that $A+K$ is diagonal, then the self-adjoint operator $A$ is said to be $\|\cdot\|_{1}$-diagonalizable.
Note that given a $\|\cdot\|_{1}$-diagonalizable self-adjoint operator $A$ and $\varepsilon>0$, there is a self-adjoint trace-class operator $K$ satisfying $\|K\|_{1}<\varepsilon$ such that $A+K$ is diagonal and each eigenvalue of $A+K$ is of multiplicity $1$. That is to say, there is an orthonormal basis $\{e_{j}\}^{\infty}_{j=1}$ of $\mathcal{H}$ such that
$A+K=\sum^{\infty}_{j=1}\alpha_{j}e_{j}\hat{\otimes}e_{j}\quad\text{and}\quad\alpha_{j}\neq\alpha_{k}\text{ for all }j\neq k.$
By the proof method adopted in [21, Halmos’ theorem] and [10, Lemma 4.33], we obtain the following result directly.
Lemma 2.1.2.
Let $T$ be an operator in $\mathcal{B}(\mathcal{H})$ with either the real part or the imaginary part being $\|\cdot\|_{1}$-diagonalizable. Then for every $\varepsilon>0$, there is a trace-class operator $K$ satisfying $\|K\|_{1}<\varepsilon$ such that $T+K$ is irreducible in $\mathcal{B}(\mathcal{H})$.
If neither the real part nor the imaginary part of $T$ is $\|\cdot\|_{1}$-diagonalizable, then we need to prepare some techniques in von Neumann algebras for later discussions.
2.2. Direct integral decomposition of operators in $\mathcal{B}(\mathcal{H})$
The theory of direct integrals is an important tool to investigate the structures of Hilbert spaces, operators, and von Neumann algebras, which is first developed by John von Neumann in [19]. While most people feel comfortable with “direct sums” of operators, many feel uneasy when they consider problems with “direct integrals” of operators. We prepare some basic knowledge of direct integrals.
Given an operator $T\in\mathcal{B}(\mathcal{H})$, we denote by $W^{*}(T)$ the von Neumann algebra generated by $T$ acting on $\mathcal{H}$. The direct integral form of $T$ depends on the choice of an abelian von Neumann subalgebra $\mathcal{D}$ in the commutant of $W^{*}(T)$. In general, $\mathcal{D}$ is not necessarily in the center of $W^{*}(T)$ (the intersection of $W^{*}(T)$ and its commutant). This is different from the type decomposition of von Neumann algebras mentioned in (1.1), where central projections play an important role.
By applying the theory of direct integrals with respect to $\mathcal{D}$, the operator $T$ can be expressed as a direct integral of operators,
$T=\int^{\oplus}_{\Lambda}T(\lambda)\mathrm{d}\mu(\lambda),\quad(\text{with respect to }\mathcal{D}).$ (2.2)
Generally, when no confusion can arise, the part “(with respect to $\mathcal{D}$)” will be omitted in (2.2). Intuitively, the larger $\mathcal{D}$ is (in the commutant of $W^{*}(T)$), the simpler the direct integrands $\{T(\lambda)\}$ become.
We briefly recall the related concepts about “direct integrals of Hilbert spaces and operators”. The reader is referred to [1, 22] for the concepts mentioned here (see Definition 4 and Definition 5 of [26, Chapter 1, Section 2] and Chapter 14 of [14] for general knowledge of direct integral of von Neumann algebras).
Fix a sequence of Hilbert spaces $\mathcal{H}_{1}\subset\mathcal{H}_{2}\subset\cdots\subset\mathcal{H}_{\infty}$ with $\mathcal{H}_{n}$ having dimension $n$ and $\mathcal{H}_{\infty}$ spanned by the union of the Hilbert spaces $\mathcal{H}_{n}$, for $1\leq n<\infty$. Next, suppose that we have a partitioned measure space $(\Lambda,\mu,\{\Lambda_{n}\}_{1\leq n\leq\infty})$, which means that $\Lambda$ is a separable metric space, $\mu$ is (the completion of) a regular Borel measure on $\Lambda$, and $\{\Lambda_{\infty},\Lambda_{1},\Lambda_{2},\ldots\}$ is a Borel partition of $\Lambda$, where some $\Lambda_{n}$’s may be of $\mu$-measure zero. We also assume that $\mu$ is $\sigma$-finite and $\Lambda$ is $\mu$-almost $\sigma$-compact. Then the symbol
$\mathcal{H}=\int^{\oplus}_{\Lambda}\mathcal{H}(\lambda)\mathrm{d}\mu(\lambda)$ (2.3)
denotes the set of all (equivalence classes of) measurable functions $f$ from $\Lambda$ into $\mathcal{H}_{\infty}$ such that
$f(\lambda)\in\mathcal{H}(\lambda)\equiv\mathcal{H}_{n}\text{ for }\lambda\in\Lambda_{n},\ 1\leq n\leq\infty\quad\text{ and }\quad\int_{\Lambda}\|f(\lambda)\|^{2}\mathrm{d}\mu(\lambda)<\infty.$ (2.4)
In addition, for $f$ and $g$ in $\mathcal{H}$ define
$\langle f,g\rangle:=\int_{\Lambda}\langle f(\lambda),g(\lambda)\rangle\mathrm{d}\mu(\lambda).$ (2.5)
The set $\mathcal{H}$ thus defined in (2.3) is called the direct integral Hilbert space with measure $\mu$ and dimension sets $\Lambda_{n}$. The vector in $\mathcal{H}$ represented by the function $\lambda\rightarrow f(\lambda)$ is denoted by $\int^{\oplus}_{\Lambda}f(\lambda)\mathrm{d}\mu(\lambda)$. It is routine to verify that with the inner product (2.5), $\mathcal{H}$ forms a complete Hilbert space.
An operator $T$ on the direct integral Hilbert space $\mathcal{H}$ is said to be decomposable if there exists a strongly $\mu$-measurable operator-valued function $T(\cdot)$ defined on $\Lambda$ such that $T(\lambda)$ is an operator in $\mathcal{B}(\mathcal{H}(\lambda))$ and $(Tf)(\lambda)=T(\lambda)f(\lambda)$, for all $f\in\mathcal{H}$. We write $T\equiv\int^{\oplus}_{\Lambda}T(\lambda)\mathrm{d}\mu(\lambda)$ for the equivalence class corresponding to $T(\cdot)$. If $T(\lambda)$ is a scalar multiple of the identity on $\mathcal{H}(\lambda)$ for almost all $\lambda$, then $T$ is said to be diagonal with respect to the direct integral Hilbert space $\mathcal{H}$. The collection of all diagonal operators is said to be the diagonal algebra $\mathcal{D}$ of $\Lambda$, which is an abelian von Neumann algebra. A decomposition ${\mathcal{A}}\thicksim\int^{\oplus}_{\Lambda}{\mathcal{A}}(\lambda)\mathrm{d}\mu(\lambda)$ of a unital strong-operator closed algebra $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$ is said to be maximal if the corresponding diagonal algebra is maximal among the abelian von Neumann subalgebras of $\mathcal{A}^{\prime}$, the commutant of $\mathcal{A}$ in $\mathcal{B}(\mathcal{H})$. These definitions are introduced preceding [1, Theorem 3.6]. By applying Zorn’s lemma, a maximal decomposition of $\mathcal{A}$ always exists. The authors of [1] proved the following theorem.
Proposition 2.2.1 (Theorem 3.6 of [1]).
Let ${\mathcal{A}}$ be a subalgebra of ${\mathcal{B}}({\mathcal{H}})$ closed in the strong operator topology. Then the decomposition ${\mathcal{A}}\thicksim\int^{\oplus}_{\Lambda}{\mathcal{A}}(\lambda)\mathrm{d}\mu(\lambda)$ is maximal if and only if almost all the algebras $\{{\mathcal{A}}(\lambda)\}$ are irreducible.
As an application, the authors of [6] proved the following proposition.
Proposition 2.2.2 (Proposition 1.1 of [6]).
Every operator in $\mathcal{B}(\mathcal{H})$ is a direct integral of irreducible operators.
Based on the preceding preparation, given an operator $T$ in $\mathcal{B}(\mathcal{H})$, there are central projections $P_{\mathrm{I}_{f}}$, $P_{\mathrm{II}_{1}}$, and $P_{\infty}$ in $W^{*}(T)$ such that we can decompose $W^{*}(T)$ with respect to (1.2) into the form
$W^{*}(T)=W^{*}(T)P_{\mathrm{I}_{f}}\oplus W^{*}(T)P_{\mathrm{II}_{1}}\oplus W^{*}(T)P_{\infty},$ (2.6)
where $W^{*}(T)P_{\mathrm{I}_{f}}$ is a finite type I von Neumann algebra, $W^{*}(T)P_{\mathrm{II}_{1}}$ is of type $\mathrm{II}_{1}$, and $W^{*}(T)P_{\infty}$ is properly infinite. Note that some of $P_{\mathrm{I}_{f}}$, $P_{\mathrm{II}_{1}}$, and $P_{\infty}$ may vanish. Furthermore, the central projections $P_{\mathrm{I}_{f}}$, $P_{\mathrm{II}_{1}}$, and $P_{\infty}$ are included in a maximal “diagonal algebra” $\mathcal{D}$ of $W^{*}(T)^{\prime}$. If $P_{\mathrm{I}_{f}}\neq 0$, then there is a family of at most countably many, mutually orthogonal projections $\{P_{j}\}_{j\geq 1}$ in $\mathcal{D}$ with sum $P_{\mathrm{I}_{f}}$ such that on each $\operatorname{ran}P_{j}$ the direct summand $W^{*}(T)P_{j}$ is unitarily equivalent to a type $\mathrm{I}_{n_{j}}$ von Neumann algebra for some certain positive integer $n_{j}$. Without loss of generality, we assume that $TP_{j}$ is an operator in a type $\mathrm{I}_{n_{j}}$ von Neumann algebra.
On the other hand, we also notice that $W^{*}(TP_{j})=W^{*}(T)P_{j}$ for each $j\geq 1$. In Section 3, we mainly prove that for each $TP_{j}$ there is an arbitrarily small trace-class operator $K_{j}$ defined on $\operatorname{ran}P_{j}$ such that $TP_{j}+K_{j}$ is a direct sum of irreducible operators. Based on this, we achieve that the direct summand $TP_{\mathrm{I}_{f}}$ can be expressed as a direct sum of irreducible operators up to an arbitrarily small trace-class perturbation.
In later sections, we adopt the following manner to describe operators in a type $\mathrm{I}_{n}$ von Neumann algebra $\mathbb{M}_{n}(L^{\infty}(\mu))$. Also, for Proposition 2.2.1, we illustrate a maximal direct integral decomposition of $\mathbb{M}_{n}(L^{\infty}(\mu))$.
Example 2.2.3.
Assume that $\mu$ is ((((the completion of)))) a finite positive regular Borel measure supported on a compact subset $\Lambda$ of $\mathbb{R}$. For the sake of simplicity, we consider elements in $L^{\infty}(\mu)$ as multiplication operators on $L^{2}(\mu)$ and matrices in $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ as bounded linear operators on $L^{2}(\mu)^{(n)}$. In this sense, every operator $T$ in $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ is of the form
$T=\begin{pmatrix}{f_{11}}&\cdots&{f_{1n}}\\ \vdots&\ddots&\vdots\\ {f_{n1}}&\cdots&{f_{nn}}\\ \end{pmatrix}_{n\times n}\begin{matrix}L^{2}(\mu)\\ \vdots\\ L^{2}(\mu)\\ \end{matrix},$ (2.7)
where the multiplication operator $M_{f_{jk}}$ is abbreviated as $f_{jk}$ in $L^{\infty}(\mu)$ and $1\leq j,k\leq n$.
[…]
As mentioned above, $L^{\infty}(\mu)$ is a maximal abelian von Neumann subalgebra in $\mathcal{B}\big{(}L^{2}(\mu)\big{)}$. Meanwhile, $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ can be viewed as a type $\mathrm{I}_{n}$ von Neumann algebra on $L^{2}(\mu)^{(n)}$. In this setting, the center $\mathcal{Z}\big{(}\mathbb{M}_{n}(L^{\infty}(\mu))\big{)}$ of $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ is in the form
$\mathcal{Z}\big{(}\mathbb{M}_{n}(L^{\infty}(\mu))\big{)}=\left\{\overbrace{f\oplus f\oplus\cdots\oplus f}^{n-{\rm fold}}\text{ acting on }L^{2}(\mu)^{(n)}:f\in L^{\infty}(\mu)\right\},$
which equals the commutant of $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$. Thus, the direct integral decomposition of $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ with respect to $\mathcal{Z}\big{(}\mathbb{M}_{n}(L^{\infty}(\mu))\big{)}$ is maximal and it can be expressed in the form
$\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}=\int^{\oplus}_{\Lambda}\mathcal{A}(\lambda)\mathrm{d}\mu(\lambda),$
where $\mathcal{A}(\lambda)=\mathbb{M}_{n}(\mathbb{C})$ for every $\lambda$ in $\Lambda$. Furthermore, if we assume that $T$ is a single generator of $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$, i.e.,
$W^{*}(T)=\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)},$ (2.8)
then by Proposition 2.2.1 we obtain that $T(\lambda)$ is irreducible in $\mathbb{M}_{n}(\mathbb{C})$ for $\lambda$ a.e. on $\Lambda$.
Sometimes, we also write $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ as $\mathbb{M}_{n}(\mathbb{C})\overline{\otimes}L^{\infty}(\mu)$, in the form of the tensor product of von Neumann algebras.
3. $\|\cdot\|_{1}$-norm perturbation of operators in finite type $\mathrm{I}$ von Neumann algebras
In this section, we prove that, for $1\leq n<\infty$, if $T\in\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ as in the form of (2.7) is reducible on $L^{2}(\mu)^{(n)}$, then it can be written as a direct sum of irreducible operators up to an arbitrarily small $\|\cdot\|_{1}$-norm perturbation.
Recall that for an operator $T$ in $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ with the assumption (2.8) in Example 2.2.3, $T(\lambda)\in\mathcal{B}(\mathcal{H}_{n})$ is irreducible a.e. on $\Lambda$, for $1\leq n<\infty$. Thus it follows that
  1. (1)
    for $n=1$, the operator $T$ is normal and by applying Theorem $4.1$ of [28], every normal operator can be written as an irreducible operator plus an arbitrarily small $\|\cdot\|_{1}$-norm perturbation;
  2. (2)
    for $2\leq n<\infty$, since $T(\lambda)\in\mathcal{B}(\mathcal{H}_{n})$ is irreducible a.e. on $\Lambda$, neither the real part nor the imaginary part of $T(\lambda)$ are scalar multiples of the identity $I_{n}$ of $\mathcal{B}(\mathcal{H}_{n})$ a.e. on $\Lambda$;
  3. (3)
    by applying [12, Theorem 3.19] to $\mathrm{Re}(T)$, the real part of $T$, there exists a unitary operator $U$ in $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ ($n\times n$ matrices with entries in $L^{\infty}(\mu)$) such that the operator $U\mathrm{Re}(T)U^{*}$ has all its non-zero entries on the diagonal of its $n\times n$ matrix form.
The following technical lemmas will be used frequently.
3.1. Preliminary lemmas
Recall that by $\mathcal{H}$ we denote a complex separable infinite-dimensional Hilbert space. For an operator $X$ in $\mathcal{B}(\mathcal{H})$, we denote by $\operatorname{ran}X$ or $X\mathcal{H}$ the range of $X$.
In the following Lemma 3.1.1, we prepare a routine construction. This lemma will be directly applied in Lemma 3.1.3 and Proposition 5.2.2. For an operator $T$ in $\mathcal{B}(\mathcal{H})$, we denote by $\sigma_{p}(T)$ the point spectrum of $T$, i.e., the set of eigenvalues of $T$. For the sake of simplicity, in a von Neumann algebra $\mathcal{M}$, a maximal abelian von Neumann subalgebra $\mathcal{A}$ is always abbreviated as a masa in $\mathcal{M}$.
Lemma 3.1.1.
Let $P$ be a nonzero projection on $\mathcal{H}$, $D$ a diagonal operator on $\operatorname{ran}P$, and $\Sigma$ a countable subset of $\mathbb{R}$. Then for every $\varepsilon>0$, there is a self-adjoint trace-class operator $K$ in $\mathcal{B}(\mathcal{H})$ of the form
$K:=\begin{pmatrix}K_{P}&0\\ 0&0\end{pmatrix}\begin{array}[]{l}\operatorname{ran}P\\ \operatorname{ran}(I-P)\end{array}$
such that
  1. $(1)$
    $\|K\|_{1}\leq\varepsilon$,
  2. $(2)$
    $\ker K_{P}=\{0\}$ i.e., $\ker K=\operatorname{ran}(I-P)$,
  3. $(3)$
    $\sigma_{p}\big{(}(D+K_{P})|_{\operatorname{ran}P}\big{)}\cap\Sigma=\varnothing$,
  4. $(4)$
    $W^{*}(D+K_{P})$ is a masa on $\operatorname{ran}P$ which is generated by minimal projections.
Proof.
Since $D$ is diagonal, there is an orthonormal basis $\{e_{j}\}_{j=1}^{N}$ for $\operatorname{ran}P$ such that $D$ is in the form $D=\sum^{N}_{j=1}\alpha_{j}e_{j}\hat{\otimes}e_{j}$, where $N$ may be infinite.
Choose a sequence $\{\delta_{j}\}^{N}_{j=1}$ of positive numbers such that for each $j$, we have that
  1. $(1)$
    $\delta_{j}\in(0,\frac{\varepsilon}{2^{j}})$,
  2. $(2)$
    $\alpha_{j}+\delta_{j}\notin\Sigma\cup\{0\}$,
  3. $(3)$
    $\alpha_{j}+\delta_{j}\neq\alpha_{k}+\delta_{k}$ for each $k=1,\ldots,j-1$.
Define $K_{P}:=\sum^{N}_{j=1}\delta_{j}e_{j}\hat{\otimes}e_{j}$ and we have that $K_{P}$ is a self-adjoint trace-class operator with $\|K_{P}\|_{1}<\varepsilon$. Thus $D+K_{P}$ is in the form
$D+K_{P}=\sum^{N}_{j=1}(\alpha_{j}+\delta_{j})e_{j}\hat{\otimes}e_{j},$
where
  1. $(1)$
    $\alpha_{j}+\delta_{j}\neq\alpha_{k}+\delta_{k}$ for all $j\neq k$ and
  2. $(2)$
    $\{\alpha_{j}+\delta_{j}\}^{N}_{j=1}\cap\Sigma=\varnothing$.
It follows that each $e_{j}\hat{\otimes}e_{j}$ is in $W^{*}(D+K_{P})$ on $\operatorname{ran}P$. Clearly, $K$ is a desired operator. This completes the proof. ∎
Remark 3.1.2.
Based on Lemma 3.1.1, we make an observation. Suppose that $A\in\mathcal{B}(\mathcal{H})$ is a self-adjoint operator in the form
$A:=\begin{pmatrix}D&0\\ 0&A_{2}\end{pmatrix}\begin{array}[]{l}\operatorname{ran}P\\ \operatorname{ran}(I-P)\end{array}$
and $\Sigma=\sigma_{p}(A_{2})$, where $D$ defined on $\operatorname{ran}P$ is diagonal. By the Borel function calculus, we have that $P\in W^{*}(A+K)$, where $K$ is given by Lemma 3.1.1.
With Lemma 3.1.1, we can perturb a class of operators $A+iB$ to be irreducible with an arbitrarily small trace-class operators.
Lemma 3.1.3.
Let $A$ and $B$ be two self-adjoint operators in $\mathcal{B}(\mathcal{H})$. If $W^{*}(A)$ contains an infinite dimensional projection $P$ with $PB=BP$ such that $(A+iB)P$ is irreducible on $P\mathcal{H}$, then for every $\varepsilon>0$, there exists a self-adjoint trace-class operator $K$ with $\|K\|_{1}\leq\varepsilon$ such that $A+i(B+K)$ is irreducible on $\mathcal{H}$.
Proof.
If $P=I$, then by the irreduciblility of $(A+iB)P$ on $P\mathcal{H}$ we choose $K=0$. In the remainder of the proof, we assume that $0<P<I$.
Since $P$ is an infinite projection, there is an isometry $V\in\mathcal{B}(\mathcal{H})$ from $(I-P)\mathcal{H}$ into $P\mathcal{H}$. Precisely, if $I-P$ is finite dimensional, then there is an isometry $V\in\mathcal{B}(\mathcal{H})$ such that
$V^{*}V=I-P\quad\text{ and }\quad VV^{*}<P.$
If $I-P$ is infinite, then we can choose $V\in\mathcal{B}(\mathcal{H})$ such that
$V^{*}V=I-P\quad\text{ and }\quad VV^{*}=P.$
By Lemma 3.1.1, there is a self-adjoint trace-class operator $K_{0}$ such that $\|K_{0}\|_{1}\leq{\varepsilon}/{2}$ and $\ker(K_{0})=\operatorname{ran}(I-P)$. Define $K:=K_{0}V+V^{*}K_{0}$. That $VP=0$ yields that $PKP=0$.
Suppose that $Q$ is a projection commuting with $A+i(B+K)$. It follows that $Q$ commutes with both $A$ and $B+K$. We will show that $Q=0$ or $I$.
By $QA=AQ$ and $P\in W^{*}(A)$, we have that $Q$ commutes with $P$. Thus, we can define two projections $Q_{1}=QP$ and $Q_{2}=Q(I-P)$. Note that
$QP(B+K)P=P(B+K)PQ.$ (3.1)
Since $PKP=0$, the equality (3.1) implies that $Q_{1}B=QPBP=PBPQ=BQ_{1}$. It follows that $Q_{1}$ is a projection commuting with $(A+iB)P$. By the hypothesis that $(A+iB)P$ is irreducible on $P\mathcal{H}$, we have $Q_{1}=0$ or $P$. Without loss of generality, we assume $Q_{1}=0$, otherwise we consider $I-Q$. By the equality $PB=BP$ and the definition of $K$, we can write $B+K$ in the form
$\begin{array}[]{rcll}B+K=&&BP&+\leavevmode\nobreak\ K_{0}V\\ &+&V^{*}K_{0}&+\leavevmode\nobreak\ B(I-P)\\ \end{array}.$
With $Q_{1}=0$ and $Q(B+K)=(B+K)Q$, we have $Q_{2}(B+K)=(B+K)Q_{2}$ and $K_{0}VQ_{2}=0$. Note that $\operatorname{ker}K_{0}=\operatorname{ran}(I-P)$ and $\operatorname{ran}V\subseteq P\mathcal{H}$. It follows that $Q_{2}=0$. Thus, $Q=0$. Therefore, $A+i(B+K)$ is irreducible. This completes the proof. ∎
In Lemma 3.1.3, if the projection $I-P$ is also infinite-dimensional, then we can remove the condition $(A+iB)P$ being irreducible on $P\mathcal{H}$, and obtain the following lemma.
Lemma 3.1.4.
Let $A$ and $B$ be two self-adjoint operators in $\mathcal{B}(\mathcal{H})$. If $W^{\ast}(A)$ contains two infinite dimensional projections $P_{1},P_{2}$ with
$P_{1}+P_{2}=I\quad\mbox{ and }\quad P_{j}B=BP_{j},\quad\mbox{ for }\ j=1,2,$ (3.2)
then for every $\varepsilon>0$, there exists a self-adjoint trace-class operator $K$ in $\mathcal{B}(\mathcal{H})$ with $\|K\|_{1}\leq\varepsilon$ such that $A+i(B+K)$ is irreducible on $\mathcal{H}$.
Proof.
The proof will be finished in two steps. First, we construct a self-adjoint trace-class operator $K$ in $\mathcal{B}(\mathcal{H})$. Then we prove that $A+i(B+K)$ is irreducible in $\mathcal{B}(\mathcal{H})$ as required.
Step 1. By the hypothesis that $P_{1}$ and $P_{2}$ are two infinite (dimensional) projections in $\mathcal{B}(\mathcal{H})$ with $P_{1}+P_{2}=I$, there exists a system of matrix units $\{P_{jk}\}^{2}_{j,k=1}$ for $\mathcal{B}(\mathcal{H})$ such that $P_{jj}=P_{j}$, for $j=1,2$. Furthermore, the projection $P_{11}$ can be decomposed into an infinite sum of mutually orthogonal rank-one projections $\{E_{jj}\}^{\infty}_{j=1}$, i.e.,
[…]
where by “SOT” we mean that the summation is convergent in the strong-operator topology. Thus, there exists a system of matrix units $\{E_{jk}\}^{\infty}_{j,k=1}$ for $P_{11}\mathcal{B}(\mathcal{H})P_{11}$. Note that the von Neumann algebra generated by $\{E_{jk}\}^{\infty}_{j,k=1}\cup\{P_{jk}\}^{2}_{j,k=1}$ is $\mathcal{B}(\mathcal{H})$ and
$I=\text{\tiny SOT-}\sum\nolimits^{\infty}_{j=1}E_{jj}+\text{\tiny SOT-}\sum\nolimits^{\infty}_{j=1}P_{21}E_{jj}P_{12}.$
To construct the imaginary part of the irreducible operator, we define two self-adjoint trace-class operators $K_{1}$ and $K_{2}$. Define $K_{1}$ in $\mathcal{B}(\mathcal{H})$ of the form
$K_{1}:=$ $\frac{\varepsilon}{2}\sum\nolimits^{\infty}_{j=1}\frac{1}{3^{j}}(E_{jj}P_{12}+P_{21}E_{jj}).$
It follows that $\|K_{1}\|_{1}\leq\varepsilon/2$.
Note that each $E_{jj}BE_{kk}$ is a scalar multiple of $E_{jk}$. We define a sequence $\{\delta_{j}\}_{j\geq 1}$ of non-negative numbers in the form
$\delta_{j}:=\begin{cases}1,&\text{if }E_{jj}BE_{j+1,j+1}=0;\\ 0,&\text{if }E_{jj}BE_{j+1,j+1}\neq 0.\\ \end{cases}$
By applying $\{\delta_{j}\}_{j\geq 1}$, we define $K_{2}$ in the form
$K_{2}:=\frac{\varepsilon}{2}\sum\nolimits^{\infty}_{j=1}\frac{\delta_{j}}{3^{j}}(E_{j,j+1}+E_{j+1,j}).$ (3.4)
It follows that $\|K_{2}\|_{1}\leq\varepsilon/2$.
Define $K:=K_{1}+K_{2}$ and $Y:=A+i(B+K)$. Thus, we have that
$\|K\|_{1}=\|K_{1}+K_{2}\|_{1}\leq\varepsilon.$
Step 2. Notice that $A$ and $B+K$ are self-adjoint operators in $W^{\ast}(Y)$, where $W^{\ast}(Y)$ is the von Neumann subalgebra of $\mathcal{B}(\mathcal{H})$ generated by $Y$.
By the hypothesis, we have
$\{P_{11},P_{22}\}\subset W^{\ast}(A)\subset W^{\ast}(Y).$
Since $P_{11}E_{jk}=E_{jk}P_{11}$ and $P_{11}P_{22}=0$, we have $P_{11}K_{2}P_{22}=0$. Moreover, by $P_{jj}B=BP_{jj}$ for $j=1,2$, we have that
$P_{11}(B+K)P_{22}=P_{11}K_{1}P_{22}=\frac{\varepsilon}{2}\sum\nolimits^{\infty}_{j=1}\frac{1}{3^{j}}(E_{jj}P_{12})\in W^{\ast}(Y).$
This implies that
$P_{11}K_{1}P_{22}K_{1}P_{11}=\frac{\varepsilon^{2}}{4}\sum\nolimits^{\infty}_{j=1}\frac{1}{3^{2j}}E_{jj}\in W^{\ast}(Y).$
By means of the function calculus for the positive operator $P_{11}K_{1}P_{22}K_{1}P_{11}$, we obtain that
$\{E_{jj}\}^{\infty}_{j=1}\subset W^{\ast}(Y).$ (3.5)
Thus for each $j\geq 1$, we have
$E_{jj}P_{12}=\frac{2\cdot 3^{j}}{\varepsilon}E_{jj}(B+K)P_{22}\in W^{\ast}(Y).$
It follows that
$P_{12}=\text{\tiny SOT-}\sum\nolimits^{\infty}_{j=1}E_{jj}P_{12}\in W^{\ast}(Y).$
Since $W^{\ast}(Y)$ is self-adjoint, we obtain that $P_{21}\in W^{*}(Y)$. Therefore, we have
$\{P_{jk}\}^{2}_{j,k=1}\subset W^{\ast}(Y).$ (3.6)
Note that $E_{jj}K_{1}E_{j+1,j+1}=0$ for all $j\geq 1$. By the definition of $K_{2}$ in (3.4) and the inclusion $\{E_{jj}\}^{\infty}_{j=1}\subseteq W^{\ast}(Y)$ in (3.5), we have that $E_{jj}(B+K)E_{j+1,j+1}\neq 0$ and
$\lambda_{j}E_{j,j+1}=\frac{E_{jj}(B+K)E_{j+1,j+1}}{\|E_{jj}(B+K)E_{j+1,j+1}\|}\in W^{\ast}(Y),\quad\forall\ j\geq 1,$ (3.7)
where $\lambda_{j}$ is some certain complex number on the unit circle. By (3.7) and that $W^{\ast}(Y)$ is self-adjoint, we have
$\{E_{jk}\}^{\infty}_{j,k=1}\subseteq W^{\ast}(Y).$ (3.8)
Note that $\{E_{jk}\}^{\infty}_{j,k=1}$ and $\{P_{jk}\}^{2}_{j,k=1}$ generate $\mathcal{B}(\mathcal{H})$ as a von Neumann algebra. It follows that $W^{\ast}(Y)=\mathcal{B}(\mathcal{H})$ by the inclusions (3.8) and (3.6). This means that $Y$ is a single generator for $\mathcal{B}(\mathcal{H})$. Thus $Y$ is irreducible in $\mathcal{B}(\mathcal{H})$. This completes the proof. ∎
To reveal a tip of the efficiency of Lemma 3.1.4, we provide a short proof of Theorem 4.1 of [28].
Corollary 3.1.5.
For each normal operator $N$ in $\mathcal{B}(\mathcal{H})$ and every $\varepsilon>0$, there is a trace-class operator $K$ in $\mathcal{B}(\mathcal{H})$ with $\|K\|_{1}<\varepsilon$ such that $N+K$ is irreducible in $\mathcal{B}(\mathcal{H})$.
Proof.
Write $N=A+iB$, where $A$ and $B$ are self-adjoint operators in $\mathcal{B}(\mathcal{H})$. Evidently, $W^{*}(A)$ is an abelian von Neumann algebra.
If $W^{*}(A)$ is finite-dimensional, then we can assume that $A$ is diagonal. We finish the proof by applying Lemma 2.1.2.
If $W^{*}(A)$ is infinite-dimensional, then there exists a sequence $\{E_{n}\}_{n=1}^{\infty}$ of nonzero projections in $W^{*}(A)$ such that $I=\text{\tiny SOT-}\sum_{n=1}^{\infty}E_{n}$. Write $P:=\text{\tiny SOT-}\sum_{n=1}^{\infty}E_{2n}$. It follows that $P$ and $I-P$ are both infinite projections in $\mathcal{B}(\mathcal{H})$. Thus the proof is completed by applying Lemma 3.1.4. ∎
Note that the operator $A+iB$ is normal if and only if $W^{*}(A)\subseteq W^{*}(A+iB)^{\prime}$. Thus, it is natural to consider Problem A for operators $A+iB$ satisfying the reverse inclusion $W^{*}(A+iB)^{\prime}\subseteq W^{*}(A)$. Before proceeding to the following Proposition 3.1.7, we make an observation in Remark 3.1.6.
Remark 3.1.6.
For every self-adjoint operators $A$ and $B$ in $\mathcal{B}(\mathcal{H})$, it is obvious to have the inclusion $W^{*}(A+iB)^{\prime}\subseteq W^{*}(A)^{\prime}$. Moreover, assume that $W^{*}(A)$ is a maximal abelian von Neumann subalgebra of $\mathcal{B}(\mathcal{H})$, which is equivalent to the inclusion $W^{*}(A)^{\prime}\subseteq W^{*}(A)$. The two inclusions imply that
$W^{*}(A+iB)^{\prime}\subseteq W^{*}(A).$ (3.9)
Generally speaking, besides the set of operators $A+iB$ with $W^{*}(A)$ generating a maximal abelian von Neumann algebra, there also exists a large family of operators satisfying (3.9), such as irreducible operators.
As an application of Lemma 3.1.4, we obtain the following proposition, which will be an important tool in the proof of Lemma 3.2.4.
Proposition 3.1.7.
Let $A$ and $B$ be two self-adjoint operators in $\mathcal{B}(\mathcal{H})$ such that
$W^{*}(A+iB)^{\prime}\subseteq W^{*}(A).$
Then for every $\varepsilon>0$, there exists a self-adjoint trace-class operator $K$ with $\|K\|_{1}\leq\varepsilon$ such that $A+i(B+K)$ is irreducible.
Proof.
Since $W^{*}(A)$ is an abelian von Neumann algebra, the hypothesis entails that $W^{*}(A+iB)^{\prime}$ is also an abelian von Neumann algebra.
If $W^{*}(A+iB)^{\prime}$ is finite-dimensional, then there is an infinite projection $P$ such that $P$ is minimal in $W^{*}(A+iB)^{\prime}$. It follows that $(A+iB)P$ is irreducible on $P\mathcal{H}$. Thus, by applying Lemma 3.1.3, there exists a self-adjoint trace-class operator $K$ with $\|K\|_{1}\leq\varepsilon$ such that $A+i(B+K)$ is irreducible.
If $W^{*}(A+iB)^{\prime}$ is infinite-dimensional, then there exists a sequence $\{E_{n}\}_{n=1}^{\infty}$ of nonzero projections in $W^{*}(A+iB)^{\prime}$ such that $I=\text{\tiny SOT-}\sum_{n=1}^{\infty}E_{n}$. Write $P:=\text{\tiny SOT-}\sum_{n=1}^{\infty}E_{2n}$. It follows that $P$ and $I-P$ are both infinite projections in $\mathcal{B}(\mathcal{H})$. Thus by applying Lemma 3.1.4. there exists a self-adjoint trace-class operator $K$ with $\|K\|_{1}\leq\varepsilon$ such that $A+i(B+K)$ is irreducible.
The above two cases complete the proof. ∎
By Proposition 3.1.7, we have a direct corollary.
Corollary 3.1.8.
Let $A$ and $B$ be two self-adjoint operators in $\mathcal{B}(\mathcal{H})$ such that $W^{*}(A)$ is a maximal abelian von Neumann subalgebra of $\mathcal{B}(\mathcal{H})$.
Then for every $\varepsilon>0$, there exists an irreducible operator $Y$ in $\mathcal{B}(\mathcal{H})$ such that
$\|(A+iB)-Y\|_{1}<\varepsilon.$
Remark 3.1.9.
One may think that for each self-adjoint operator $A$ in $\mathcal{B}(\mathcal{H})$, if there exists an arbitrarily small self-adjoint trace-class operator $K$ such that $W^{*}(A+K)$ is a masa in $\mathcal{B}(\mathcal{H})$, then Problem A can be solved completely by applying Corollary 3.1.8. But the thought fails to work. We provide such a self-adjoint operator without a proof. Let $M_{t}$ be the multiplication operator on $L^{2}[0,1]$ defined by $(M_{t}f)(t):=t\cdot f(t)$ for every $t\in[0,1]$ and every $f\in L^{2}[0,1]$. Let $\mathcal{H}=L^{2}[0,1]\oplus L^{2}[0,1]$ and $A:=M_{t}\oplus M_{t}$. Clearly, $W^{*}(A)$ is not a masa in $\mathcal{B}(\mathcal{H})$. By Theorem $5.2.5$ of [17], for each self-adjoint trace-class operator $K$, $W^{*}(A+K)$ also fails to be a masa in $\mathcal{B}(\mathcal{H})$.
3.2. $\|\cdot\|_{1}$-norm perturbation of operators in type $\mathrm{I}_{n}$ von Neumann algebras
Based on the preparation in Subsection 3.1, we start by considering operators in $\mathbb{M}_{2}\big{(}L^{\infty}(\mu)\big{)}$. For every operator $T\in\mathbb{M}_{2}\big{(}L^{\infty}(\mu)\big{)}$, we prove that $T$ can be expressed as a direct sum of irreducible operators up to an arbitrarily small trace-class perturbation. Then we generalize this result for $T_{n}\in\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ and $\bigoplus^{\infty}_{n=1}T_{n}$. A key tool is Lemma 3.1.4 and we need to discuss the spectra of the diagonal entries of the real part of $T_{n}$.
To prove the result for $T$ in $\mathbb{M}_{2}\big{(}L^{\infty}(\mu)\big{)}$, we start by considering a spacial case for $T$ in Lemma 3.2.3. Then by Lemma 3.2.3, we prove Proposition 3.2.1. By a series of similar arguments, we prove the result for $T_{n}\in\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ in Proposition 3.2.6. In Proposition 3.2.7, we prove the result for $\bigoplus^{\infty}_{n=1}T_{n}$.
In this subsection, we assume that $\mu$ is (the completion of) a finite positive regular Borel measure supported on a $\sigma$-compact subset $\Lambda$ of $\mathbb{R}$. Write $L^{\infty}(\mu)$ to be the set of equivalence classes of essentially bounded $\mu$-measurable functions on $\Lambda$, where the equivalence relation $f\sim g$ holds if and only if $f=g$ a.e. on $\Lambda$. For the sake of simplicity of language, we shall follow the convention to view $L^{\infty}(\mu)$ as a space of functions (see [13, Example 1.7.4] for more details).
Meanwhile, there exists a cyclic faithful normal $\ast$-representation $\pi:L^{\infty}(\mu)\to\mathcal{B}\big{(}L^{2}(\mu)\big{)}$ defined by $\pi(\phi):=M_{\phi}$, where by $M_{\phi}$ we denote the multiplication operator with symbol $\phi$. Write
$\mathcal{A}_{\mu}:=\{M_{\phi}:\phi\in L^{\infty}(\mu)\}.$ (3.10)
Thus, $\pi$ is a normal $\ast$-isomorphism from $L^{\infty}(\mu)$ onto $\mathcal{A}_{\mu}$. It is routine to check $\mathcal{A}_{\mu}$ is a masa in $\mathcal{B}\big{(}L^{2}(\mu)\big{)}$.
Recall that, for every $\phi$ in $L^{\infty}(\mu)$, the essential range of $\phi$ (see [13, Example 3.2.16] or [4, Definition 2.61]) equals the spectrum $\sigma(M_{\phi})$ of the multiplication operator $M_{\phi}$ acting on $L^{2}(\mu)$. Since $\pi$ in the preceding paragraph is a $\ast$-isomorphism, we have $\sigma(\phi)=\sigma(M_{\phi})$. It follows that $\sigma(\phi)$ equals the essential range of $\phi$.
In the following proposition, we consider self-adjoint operators $A$ and $B$ in the type $\mathrm{I}_{2}$ von Neumann algebra ${\mathbb{M}}_{2}\big{(}L^{\infty}(\mu)\big{)}$ acting on $L^{2}(\mu)^{(2)}$ satisfying that $(A+iB)(\lambda)$ is in ${\mathbb{M}}_{2}(\mathbb{C})$ for every $\lambda$ in $\Lambda$. We construct an arbitrarily small trace-class perturbation $K$ such that $(A+iB)+K$ is a direct sum of at most countably many irreducible operators.
For the sake of simplicity, when dealing with matrices with entries from $\mathcal{A}_{\mu}$ we consider elements $\phi$ in $L^{\infty}(\mu)$ as multiplication operators $M_{\phi}$ acting on $L^{2}(\mu)$ and the cyclic, faithful, normal $\ast$-representation $\pi:L^{\infty}(\mu)\to\mathcal{B}\big{(}L^{2}(\mu)\big{)}$ defined by $\pi(\phi):=M_{\phi}$ preceeding (3.10) will be omitted when no confusion can arise.
Proposition 3.2.1.
Suppose that $A$ and $B$ are two self-adjoint operators in ${\mathbb{M}}_{2}\big{(}L^{\infty}(\mu)\big{)}$ acting on $L^{2}(\mu)^{(2)}$ defined as follows,
$A:=\begin{pmatrix}{f}&0\\ 0&g\\ \end{pmatrix}\begin{matrix}L^{2}(\mu)\\ L^{2}(\mu)\\ \end{matrix},\quad B:=\begin{pmatrix}{h_{11}}&h_{12}\\ h_{21}&h_{22}\\ \end{pmatrix}\begin{matrix}L^{2}(\mu)\\ L^{2}(\mu)\\ \end{matrix},\quad\{f,g,h_{jk}:1\leq j,k\leq 2\}\subset L^{\infty}(\mu).$ (3.11)
Then for every $\varepsilon>0$, there is a trace-class operator $K$ satisfying $\|K\|_{1}<\varepsilon$ such that $(A+iB)+K$ is a direct sum of at most countably many irreducible operators in $\mathcal{B}\big{(}L^{2}(\mu)^{(2)}\big{)}$.
To prove Proposition 3.2.1, we prepare a technical lemma to construct a $\mu$-measurable, at most countable partition $\{\Lambda_{k}\}_{k=1}^{N}$ with a nice property. Notice that a real-valued $\mu$-measurable function $f$ is finite for almost every $\lambda$ in $\Lambda$.
Lemma 3.2.2.
Suppose that $(\Lambda,\mu)$ is a measure space. If $f_{1},f_{2},\ldots,f_{n}$ are real-valued $\mu$-measurable functions on $\Lambda$ such that $f_{j}(\lambda)\neq f_{k}(\lambda)$ for every $\lambda\in\Lambda$ and all $j\neq k$. Then there is an at most countable $\mu$-measurable partition $\{\Lambda_{\ell}\}_{\ell=1}^{N}$ of $\Lambda$ such that
$\overline{f_{j}(\Lambda_{\ell})}\cap\overline{f_{k}(\Lambda_{\ell})}=\varnothing,\quad\text{ for all }\ell\geq 1\text{ and }j\neq k,$
where $N$ may take $\infty$.
Proof.
Let $\{G_{m}\}_{m=1}^{\infty}$ be the family of all closed intervals $[a,b]$ with rational endpoints $a,b\in\mathbb{Q}$. Let $\{(m_{1,\ell},m_{2,\ell},\ldots,m_{n,\ell})\}_{\ell=1}^{\infty}$ be the set
$\{(m_{1,\ell},m_{2,\ell},\ldots,m_{n,\ell})\}_{\ell=1}^{\infty}:=\{(m_{1},m_{2},\ldots,m_{n})\in\mathbb{N}^{(n)}\colon G_{m_{j}}\cap G_{m_{k}}=\varnothing\leavevmode\nobreak\ \text{ for }\leavevmode\nobreak\ j\neq k\}.$
For each positive integer $\ell$, we define a measurable set $\Omega_{\ell}\subseteq\Lambda$ by
$\Omega_{\ell}:=f_{1}^{-1}(G_{m_{1,\ell}})\cap f_{2}^{-1}(G_{m_{2,\ell}})\cap\cdots\cap f_{n}^{-1}(G_{m_{n,\ell}}).$
It is clear that
$\overline{f_{j}(\Omega_{\ell})}\cap\overline{f_{k}(\Omega_{\ell})}\subseteq G_{m_{j,\ell}}\cap G_{m_{k,\ell}}=\varnothing.$
By the hypothesis, we can assume that for every $\lambda\in\Lambda$ and each $j=1,\ldots,n$, there are disjoint intervals $G_{m_{1}},G_{m_{2}},\ldots,G_{m_{n}}$ such that $f_{j}(\lambda)\in G_{m_{j}}$. Then there exists some $\ell$ such that
$(m_{1,\ell},m_{2,\ell},\ldots,m_{n,\ell})=(m_{1},m_{2},\ldots,m_{n}).$
It follows that $\lambda\in\Omega_{\ell}$. Thus $\Lambda=\bigcup_{\ell=1}^{\infty}\Omega_{\ell}$. We define
$\Lambda_{1}:=\Omega_{1},\quad\Lambda_{2}:=\Omega_{2}\setminus\Omega_{1},\quad\Lambda_{3}:=\Omega_{3}\setminus(\Omega_{1}\cup\Omega_{2}),\quad\ldots.$
Therefore, we obtain an at most countable $\mu$-measurable partition $\{\Lambda_{\ell}\}_{\ell=1}^{N}$ of $\Lambda$ with the desired properties, where $N$ may be $\infty$. ∎
Note that the partition $\{\Lambda_{\ell}\}_{\ell=1}^{N}$ of $\Lambda$ obtained in Lemma 3.2.2 is not unique in general. By taking different orderings of $\{G_{m}\}_{m=1}^{\infty}$, we obtain different partitions $\{\Lambda_{\ell}\}_{\ell=1}^{N}$. Also, we should notice that the partition is possibly finite.
Recall that in $\mathcal{B}(\mathcal{H})$ the trace-class norm is stronger than the operator norm, i.e., for each $T$ in $\mathcal{B}(\mathcal{H})$, $\|T\|\leq\|T\|_{1}$. For two operators $T_{1}$ and $T_{2}$ in $\mathcal{B}(\mathcal{H})$, suppose that $\sigma(T_{1})\cap\sigma(T_{2})=\varnothing$. Then by the upper semi-continuity of the spectrum (see [25, Theorem 10.20]), there is a $\delta>0$ such that for every $X_{1}$ and $X_{2}$ with $\max\{\|X_{1}\|_{1},\|X_{2}\|_{1}\}<\delta$, we have
$\sigma(T_{1}+X_{1})\cap\sigma(T_{2}+X_{2})=\varnothing.$
In the following lemma, we prove a special case of Proposition 3.2.1.
Lemma 3.2.3.
Suppose that $A$ and $B$ are two self-adjoint operators in $\mathbb{M}_{2}\big{(}L^{\infty}(\mu)\big{)}$ acting on $L^{2}(\mu)^{(2)}$ defined as in (3.11) such that $\sigma(f)\cap\sigma(g)=\varnothing$.
Then for every $\varepsilon>0$, there is a trace-class operator $K$ satisfying $\|K\|_{1}<\varepsilon$ such that $(A+iB)+K$ is an irreducible operator in $\mathcal{B}\big{(}L^{2}(\mu)^{(2)}\big{)}$.
Proof.
As shown in (3.11), $f+ih_{11}$ and $g+ih_{22}$ are both normal operators in $\mathcal{B}\big{(}L^{2}(\Lambda,\mu)\big{)}$. If $h_{12}=0$, then $h_{21}=0$ and the operator $A+iB$ is normal. By applying Corollary 3.1.5 directly, there is a trace-class operator $K$ in $\mathcal{B}\big{(}L^{2}(\mu)^{(2)}\big{)}$ with $\|K\|_{1}<\varepsilon$ such that $(A+iB)+K$ is irreducible in $\mathcal{B}\big{(}L^{2}(\mu)^{(2)}\big{)}$.
In the remainder of the proof, we assume that $h_{12}\neq 0$. By Corollary 3.1.5 and [25, Theorem 10.20] (see the argument preceding this lemma), there are trace-class operators $K_{j}$ in $\mathcal{B}\big{(}L^{2}(\Lambda,\mu)\big{)}$ with $\|K_{j}\|_{1}<\varepsilon/2$ for $j=1,2$ such that
  1. $(1)$
    $(f+ih_{11})+K_{1}$ and $(g+ih_{22})+K_{2}$ are both irreducible in $\mathcal{B}\big{(}L^{2}(\Lambda,\mu)\big{)}$,
  2. $(2)$
    $\sigma(f+\operatorname{Re}K_{1})\cap\sigma(g+\operatorname{Re}K_{2})=\varnothing$.
Define a trace-class operator $K$ of the form
$K:=\begin{pmatrix}K_{1}&0\\ 0&K_{2}\end{pmatrix}\begin{matrix}L^{2}(\Lambda,\mu)\\ L^{2}(\Lambda,\mu)\end{matrix}$
Thus we have $\|K\|_{1}<\varepsilon$.
By the assumption that $h_{12}\neq 0$ and applying Lemma 2.1.1, the operator $(A+iB)+K$ is irreducible in $\mathcal{B}\big{(}L^{2}(\mu)^{(2)}\big{)}$. This completes the proof. ∎
We are ready to prove Proposition 3.2.1 with applying Lemma 3.2.3 and Lemma 3.2.2.
Proof of Proposition 3.2.1.
By changing the values of $f$ and $g$ on a subset of $\mu$-measure zero, we take a Borel partition of $\Lambda$ in the form $\Lambda=\Lambda^{\prime}\cup\Lambda^{\prime\prime}$ such that
  1. $(1)$
    $f(\lambda)=g(\lambda)$ for every $\lambda$ in $\Lambda^{\prime}$,
  2. $(2)$
    $f(\lambda)\neq g(\lambda)$ for every $\lambda$ in $\Lambda^{\prime\prime}$.
It is clear that the operator $A+iB$ restricted in $\mathcal{B}\big{(}L^{2}(\Lambda^{\prime},\mu)^{(2)}\big{)}$ is normal. Thus by applying Corollary 3.1.5, there is a trace-class operator $K_{0}$ on $L^{2}(\Lambda^{\prime},\mu)^{(2)}$ with $\|K_{0}\|_{1}<\varepsilon/2$ such that the restricted operator $(A+iB)|_{L^{2}(\Lambda^{\prime},\mu)^{(2)}}+K_{0}$ is irreducible in $\mathcal{B}\big{(}L^{2}(\Lambda^{\prime},\mu)^{(2)}\big{)}$.
By Lemma 3.2.2, there is an at most countable $\mu$-measurable partition $\{\Lambda_{\ell}\}^{N}_{\ell=1}$ of $\Lambda^{\prime\prime}$ such that
$\overline{f(\Lambda_{\ell})}\cap\overline{g(\Lambda_{\ell})}=\varnothing,\quad\text{ for all }\ell\geq 1,$
where $N$ may take $\infty$.
By applying Lemma 3.2.3, for each $\ell\geq 1$, there exists a trace-class operator $K_{\ell}$ in $\mathcal{B}\big{(}L^{2}(\Lambda_{\ell},\mu)^{(2)}\big{)}$ with $\|K_{\ell}\|_{1}<\varepsilon/2^{\ell+1}$ such that the operator
$(A+iB)|_{L^{2}(\Lambda_{\ell},\mu)^{(2)}}+K_{\ell}$
on $L^{2}(\Lambda_{\ell},\mu)^{(2)}$ is irreducible.
Write $K=\bigoplus^{N}_{\ell=0}K_{\ell}$. It follows that $(A+iB)+K$ is a direct sum of at most countably many irreducible operators and $K\in\mathcal{B}\big{(}L^{2}(\Lambda,\mu)^{(2)}\big{)}$ is a trace-class operator with $\|K\|_{1}<\varepsilon$. This completes the proof. ∎
With respect to Lemma 3.2.3, our aim in this section is to prove a generalization of Proposition 3.2.1 for operators in a type $\mathrm{I}_{n}$ von Neumann algebra.
Lemma 3.2.4.
Let $r_{1},r_{2},\ldots,r_{m}$ be positive integers satisfying $r_{1}+r_{2}+\cdots+r_{m}=n$ and
$\{f_{j},h_{kl}\in L^{\infty}(\mu):f_{j}=\overline{f_{j}}\text{ for }1\leq j\leq m;h_{kl}=\overline{h_{lk}}\text{ for }1\leq k,l\leq n\}.$
Suppose that $A$ and $B$ are two self-adjoint operators in $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ acting on $L^{2}(\mu)^{(n)}$ in the form
$A:=\bigoplus^{m}_{j=1}M^{(r_{j})}_{f_{j}}\quad\text{ and }\quad B:=\begin{pmatrix}h_{11}&h_{12}&\cdots&h_{1n}\\ h_{21}&h_{22}&\cdots&h_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ h_{n1}&h_{n2}&\cdots&h_{nn}\\ \end{pmatrix}\begin{matrix}L^{2}(\mu)\\ L^{2}(\mu)\\ \vdots\\ L^{2}(\mu)\\ \end{matrix}$ (3.12)
satisfying
$\sigma(f_{j})\cap\sigma(f_{k})=\varnothing\quad\text{ for all }\ j\neq k.$
Then for every $\varepsilon>0$, there exists a trace-class operator $K$ satisfying $\|K\|_{1}<\varepsilon$ such that $(A+iB)+K$ is an irreducible operator in $\mathcal{B}\big{(}L^{2}(\mu)^{(n)}\big{)}$.
Proof.
Rewrite $A$ and $B$ in the following forms
$A=\begin{pmatrix}A_{1}&0&\cdots&0\\ 0&A_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&A_{m}\\ \end{pmatrix}\begin{matrix}L^{2}(\mu)^{(r_{1})}\\ L^{2}(\mu)^{(r_{2})}\\ \vdots\\ L^{2}(\mu)^{(r_{m})}\\ \end{matrix}\quad\text{and}\quad B=\begin{pmatrix}B_{11}&B_{12}&\cdots&B_{1m}\\ B_{21}&B_{22}&\cdots&B_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ B_{m1}&B_{m2}&\cdots&B_{mm}\\ \end{pmatrix}\begin{matrix}L^{2}(\mu)^{(r_{1})}\\ L^{2}(\mu)^{(r_{2})}\\ \vdots\\ L^{2}(\mu)^{(r_{m})}\\ \end{matrix}$
where each $A_{j}:=M^{(r_{j})}_{f_{j}}$ and $B_{jk}\in\mathbb{M}_{r_{j}\times r_{k}}\big{(}L^{\infty}(\mu)\big{)}$ for $j,k=1,\ldots,m$.
Note that the operator $A_{j}+iB_{jj}$ is normal for each $j=1,\ldots,m$. By Corollary 3.1.5 and [25, Theorem 10.20], for each normal operator $A_{j}+iB_{jj}$, there is a trace-class operator $K_{jj}$ in $\mathcal{B}\big{(}L^{2}(\Lambda,\mu)^{(r_{j})}\big{)}$ such that
  1. $(1)$
    $(A_{j}+iB_{jj})+K_{jj}$ is irreducible in $\mathcal{B}\big{(}L^{2}(\Lambda,\mu)^{(r_{j})}\big{)}$,
  2. $(2)$
    $\|K_{jj}\|_{1}<\varepsilon/(2m)$, and
  3. $(3)$
    $\sigma(A_{j}+\operatorname{Re}K_{jj})\cap\sigma(A_{k}+\operatorname{Re}K_{kk})=\varnothing$ for all $j\neq k$.
Write $K_{1}:=\bigoplus^{m}_{j=1}K_{jj}$. Then we have $\|K_{1}\|_{1}<\varepsilon/2$.
We verify that
$W^{*}((A+iB)+K_{1})^{\prime}\subseteq W^{*}(A+\operatorname{Re}K_{1}).$ (3.13)
Assume that $P$ is a projection in $\mathcal{B}\big{(}L^{2}(\mu)^{(n)}\big{)}$ commuting with $(A+iB)+K_{1}$. Since the relation $\sigma(A_{j}+\operatorname{Re}K_{jj})\cap\sigma(A_{k}+\operatorname{Re}K_{kk})=\varnothing$ holds for all $j\neq k$, the spectral projection $E_{j}$ for $A+\operatorname{Re}K_{1}$ corresponding to $\sigma(A_{j}+\operatorname{Re}K_{jj})$ is exactly the projection onto the $j$-th direct summand Hilbert space $L^{2}(\mu)^{(r_{j})}$. The equality $P(A+\operatorname{Re}K_{1})=(A+\operatorname{Re}K_{1})P$ yields $PE_{j}=E_{j}P$ for each $1\leq j\leq m$. Thus, $P$ can be expressed in the form
$P=\begin{pmatrix}P_{1}&0&\cdots&0\\ 0&P_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&P_{m}\\ \end{pmatrix}\begin{matrix}L^{2}(\mu)^{(r_{1})}\\ L^{2}(\mu)^{(r_{2})}\\ \vdots\\ L^{2}(\mu)^{(r_{m})}\\ \end{matrix}.$ (3.14)
Note that one can also deduce the matrix form of $P$ as in (3.14) by a property of Rosenblum operator (see Corollary 0.13 of [23]). Moreover, for each $1\leq j\leq m$, the equality $P_{j}(B_{jj}+\operatorname{Im}K_{jj})=(B_{jj}+\operatorname{Im}K_{jj})P_{j}$ and the irreducibility of $(A_{j}+\operatorname{Re}K_{jj})+i(B_{jj}+\operatorname{Im}K_{jj})$ yield that $P_{j}$ is a trivial projection on $L^{2}(\mu)^{(r_{j})}$. This verifies (3.13). Thus, by applying Proposition 3.1.7, there exists a self-adjoint trace-class operator $K_{2}$ with $\|K_{2}\|_{1}<\varepsilon/2$ such that $(A+iB)+K_{1}+iK_{2}$ is irreducible in $\mathcal{B}\big{(}L^{2}(\mu)^{(n)}\big{)}$. Write $K:=K_{1}+iK_{2}$. This completes the proof. ∎
Next, we generalize Lemma 3.2.4 by replacing the condition $\sigma(f_{j})\cap\sigma(f_{k})=\varnothing$ for all $j\neq k$ with a weaker one that $f_{j}(\lambda)\neq f_{k}(\lambda)$ for all $j\neq k$ and almost every $\lambda$ in $\Lambda$. The method is similar to the proof of Proposition 3.2.1 by combining Lemma 3.2.2 and Lemma 3.2.4.
Lemma 3.2.5.
Let $r_{1},r_{2},\ldots,r_{m}$ be positive integers satisfying $r_{1}+r_{2}+\cdots+r_{m}=n$ and
$\{f_{j},h_{kl}\in L^{\infty}(\mu):f_{j}=\overline{f_{j}}\text{ for }1\leq j\leq m;h_{kl}=\overline{h_{lk}}\text{ for }1\leq k,l\leq n\}.$
Suppose that $A$ and $B$ are two self-adjoint operators in $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ acting on $L^{2}(\mu)^{(n)}$ in the form
$A:=\bigoplus^{m}_{j=1}M^{(r_{j})}_{f_{j}}\quad\text{ and }\quad B:=\begin{pmatrix}h_{11}&h_{12}&\cdots&h_{1n}\\ h_{21}&h_{22}&\cdots&h_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ h_{n1}&h_{n2}&\cdots&h_{nn}\\ \end{pmatrix}\begin{matrix}L^{2}(\mu)\\ L^{2}(\mu)\\ \vdots\\ L^{2}(\mu)\\ \end{matrix}$ (3.15)
satisfying $f_{j}(\lambda)\neq f_{k}(\lambda)$ for all $j\neq k$ and almost every $\lambda$ in $\Lambda$.
Then for every $\varepsilon>0$, there exists a trace-class operator $K$ satisfying $\|K\|_{1}<\varepsilon$ such that $(A+iB)+K$ is a direct sum of at most countably many irreducible operators in $\mathcal{B}\big{(}L^{2}(\mu)^{(n)}\big{)}$.
Proof.
By changing values of $f_{j}$’s on a set of $\mu$-measure zero, we assume that $f_{j}(\lambda)\neq f_{k}(\lambda)$ for all $j\neq k$ and every $\lambda$ in $\Lambda$.
By applying Lemma 3.2.2, there is an at most countable $\mu$-measurable partition $\{\Lambda_{\ell}\}^{N}_{\ell=1}$ of $\Lambda$ such that
$\overline{f_{j}(\Lambda_{\ell})}\cap\overline{f_{k}(\Lambda_{\ell})}=\varnothing,\quad\text{ for all }\ell\geq 1\text{ and }j\neq k,$
where $N$ may take $\infty$.
On every $L^{2}(\Lambda_{\ell},\mu)^{(n)}$, applying Lemma 3.2.4, we obtain a trace-class operator $K_{\ell}$ with $\|K_{\ell}\|_{1}<\varepsilon/2^{\ell}$ such that the restriction
$(A+iB)|_{L^{2}(\Lambda_{\ell},\mu)^{(n)}}+K_{\ell}$
on $L^{2}(\Lambda_{\ell},\mu)^{(n)}$ is irreducible.
Write $K=\bigoplus^{N}_{\ell=1}K_{\ell}$. It follows that $(A+iB)+K$ is a direct sum of at most countably many irreducible operators and $K\in\mathcal{B}\big{(}L^{2}(\Lambda,\mu)^{(n)}\big{)}$ is a trace-class operator with $\|K\|_{1}<\varepsilon$. This completes the proof. ∎
Based on Lemma 3.2.5, we achieve the following proposition for $T$ in $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$.
Proposition 3.2.6.
Suppose that $A$ and $B$ are two self-adjoint operators in $\mathbb{M}_{n}\big{(}L^{\infty}(\mu)\big{)}$ acting on $L^{2}(\mu)^{(n)}$ in the form
$A:=\bigoplus^{n}_{j=1}M_{f_{j}}\quad\text{ and }\quad B:=\begin{pmatrix}h_{11}&h_{12}&\cdots&h_{1n}\\ h_{21}&h_{22}&\cdots&h_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ h_{n1}&h_{n2}&\cdots&h_{nn}\\ \end{pmatrix}\begin{matrix}L^{2}(\mu)\\ L^{2}(\mu)\\ \vdots\\ L^{2}(\mu)\\ \end{matrix}.$
Then for every $\varepsilon>0$, there is a trace-class operator $K$ satisfying $\|K\|_{1}<\varepsilon$ such that $(A+iB)+K$ is a direct sum of at most countably many irreducible operators in $\mathcal{B}\big{(}L^{2}(\mu)^{(n)}\big{)}$.
Proof.
Note that there is a finite $\mu$-measurable partition $\{\Lambda_{\ell}\}^{N}_{\ell=1}$ of $\Lambda$ such that on each $\Lambda_{\ell}$ there are positive integers $r_{\ell,1},r_{\ell,2},\ldots,r_{\ell,m_{\ell}}$ with sum $n$ and $\{f_{1},f_{2},\ldots,f_{n}\}$ can be rewritten as
$\{\overbrace{g_{\ell,1},\ldots,g_{\ell,1}}^{r_{\ell,1}-{\rm fold}},\overbrace{g_{\ell,2},\ldots,g_{\ell,2}}^{r_{\ell,2}-{\rm fold}},\ldots,\overbrace{g_{\ell,m_{\ell}},\ldots,g_{\ell,m_{\ell}}}^{r_{\ell,m_{\ell}}-{\rm fold}}\}$
satisfying $g_{\ell,j}(\lambda)\neq g_{\ell,k}(\lambda)$ for all $j\neq k$ and almost every $\lambda$ in $\Lambda_{\ell}$. By applying Lemma 3.2.5, there exists a trace-class operator $K_{\ell}$ on $L^{2}(\Lambda_{\ell},\mu)^{(n)}$ with $\|K_{\ell}\|<\varepsilon/2^{\ell}$ such that the restriction
$(A+iB)|_{L^{2}(\Lambda_{\ell},\mu)^{(n)}}+K_{\ell}$
on $L^{2}(\Lambda_{\ell},\mu)^{(n)}$ is a direct sum of at most countably many irreducible operators.
Write $K=\bigoplus^{N}_{\ell=1}K_{\ell}$. It follows that $(A+iB)+K$ is a direct sum of at most countably many irreducible operators and $K\in\mathcal{B}\big{(}L^{2}(\Lambda,\mu)^{(n)}\big{)}$ is a trace-class operator with $\|K\|_{1}<\varepsilon$. This completes the proof. ∎
Combining with the theory of direct integrals which is introduced in Subsection 2.2 and Proposition 3.2.6, we obtain the following proposition.
Proposition 3.2.7.
Suppose that $T$ is an operator in a finite type $\mathrm{I}$ von Neumann algebra. Then for every $\varepsilon>0$, there is a trace-class operator $K$ satisfying $\|K\|_{1}<\varepsilon$ such that the operator $T+K$ is a direct sum of at most countably many irreducible operators.
Proof.
Since $T$ is in a finite type I von Neumann algebra, following the notation of direct integrals as introduced in Subsection 2.2, up to unitary equivalence, we can write
$T=\bigoplus^{\infty}_{n=1}T_{n}\quad\text{ and }\quad T_{n}\in\mathbb{M}_{n}\big{(}L^{\infty}(\Lambda_{n},\mu)\big{)}\text{ acting on }L^{2}(\Lambda_{n},\mu)^{(n)},$
where some $\Lambda_{n}$’s may be of $\mu$-measure zero, and the direct sum is possibly finite. In each $\Lambda_{n}$, we apply Proposition 3.2.6 to each $T_{n}$ with respect to the subset $\Lambda_{n}$ to obtain a trace-class operator $K_{n}$ on $L^{2}(\Lambda_{n},\mu)^{(n)}$ with $\|K_{n}\|_{1}<\varepsilon/2^{n}$ such that $T_{n}+K_{n}$ is a direct sum of at most countably many irreducible operators. Write $K:=\bigoplus_{n\geq 1}K_{n}$. Therefore, we have that $\|K\|_{1}<\varepsilon$ and the operator $T+K$ is a direct sum of at most countably many irreducible operators. This completes the proof. ∎
4. Perturbation of single generators of properly infinite von Neumann algebras in trace-class norm
Recall that, in (2.6), for an operator $T$ in $\mathcal{B}(\mathcal{H})$, there are at most three direct summands $T_{1}$, $T_{2}$, and $T_{\infty}$ satisfying $T=T_{1}\oplus T_{2}\oplus T_{\infty}$, where $W^{*}(T_{1})$ is a finite type I von Neumann algebra, $W^{*}(T_{2})$ is a type $\mathrm{II}_{1}$ von Neumann algebra, and $W^{*}(T_{\infty})$ is a properly infinite von Neumann algebra. In this section, we prove that an operator as in (4.4) can be expressed as an irreducible operator up to an arbitrarily small trace-class perturbation.
Precisely, we start by proving in Lemma 4.1.2 that, for a self-adjoint operator $A$, by a perturbation of a self-adjoint trace-class operator $K$, the self-adjoint operator $A+K$ can always have an isolated eigenvalue with multiplicity $1$. Then we develop a method to perturb a class of operators constructed with $T_{\infty}$ with respect to an arbitrarily small trace-class norm.
4.1. Finite rank perturbation of self-adjoint operators
In (2.1), we recall a rank-one operator $e\hat{\otimes}f$ on $\mathcal{H}$ defined with two vectors $e$ and $f$ in $\mathcal{H}$. By $\mathrm{Tr}(\cdot)$ we denote the standard trace on the set of trace-class operators in $\mathcal{B}(\mathcal{H})$.
Remark 4.1.1.
For rank-one operators we recall the following properties. If $e$ and $f$ are two unit vectors in $\mathcal{H}$, then the rank-one operator $e\hat{\otimes}f$ is a partial isometry with initial projection $f\hat{\otimes}f$ and final projection $e\hat{\otimes}e$. In general, for two vectors $e$ and $f$ in $\mathcal{H}$, we have
$\|e\hat{\otimes}f\|_{1}=\|e\|\|f\|=\|e\hat{\otimes}f\|.$ (4.1)
It is routine to verify the equality (4.1). By definition, we have $\|e\hat{\otimes}f\|_{1}=\mathrm{Tr}(|e\hat{\otimes}f|)$. If either $e$ or $f$ is zero, then (4.1) holds automatically. If $e$ and $f$ are both nonzero, then without loss of generality we can verify (4.1) by assuming that $\|e\|=\|f\|=1$.
For every self-adjoint operator $A$, we denote by $E_{A}(\,\cdot\,)$ the spectral measure for $A$. In the following lemma, we recall a classical technique of constructing an arbitrarily small self-adjoint trace-class perturbation $K$ of $A$ such that $\sigma(A+K)$ contains an isolated eigenvalue of multiplicity $1$. Recall that we denote by $\|\cdot\|_{1}$-norm the Schatten 1-norm (in other words, trace-class norm).
Lemma 4.1.2.
Let $A$ be a self-adjoint operator in $\mathcal{B}(\mathcal{H})$. Then for every $\varepsilon>0$, there exists a self-adjoint finite rank operator $K$ such that
  1. $(1)$
    $\|K\|_{1}\leq\varepsilon$, and
  2. $(2)$
    there is an isolated eigenvalue of $A+K$ with multiplicity $1$.
Proof.
Without loss of generality, we assume that $A$ is positive and $\|A\|=1$. In this case, we have that $1\in\sigma(A)$. Define a Borel set $\Delta_{\varepsilon}$ of the form
$\Delta_{\varepsilon}:=\sigma(A)\cap[1-\frac{\varepsilon}{4},1].$
By the assumption, $E_{A}(\Delta_{\varepsilon})$ is a nonzero spectral projection for $A$ and
$\|(A-I)E_{A}(\Delta_{\varepsilon})\|\leq\frac{\varepsilon}{4}.$
Choose a unit vector $\xi$ in $\operatorname{ran}E_{A}(\Delta_{\varepsilon})$ and denote by $F_{\xi}$ the rank-one projection $\xi\hat{\otimes}\xi$ onto $\mathbb{C}\xi$. Then we have
$\|F_{\xi}(A-I)F_{\xi}\|_{1}\leq\|(A-I)F_{\xi}\|_{1}=\|(A-I)E_{A}(\Delta_{\varepsilon})F_{\xi}\|_{1}\leq\|(A-I)E_{A}(\Delta_{\varepsilon})\|\|F_{\xi}\|_{1}\leq\frac{\varepsilon}{4}.$
Since $F_{\xi}(A-I)=((A-I)F_{\xi})^{*}$, we have that $\|F_{\xi}(A-I)\|_{1}=\|(A-I)F_{\xi}\|_{1}$. Define $K$ to be in the form
$K:=\frac{\varepsilon}{4}F_{\xi}+F_{\xi}(A-I)F_{\xi}-F_{\xi}(A-I)-(A-I)F_{\xi}.$
Then $K$ is of finite rank with $\|K\|_{1}\leq\varepsilon$ and $A+K$ is in the form
$A+K=(1+\frac{\varepsilon}{4})F_{\xi}+(I-F_{\xi})A(I-F_{\xi}).$ (4.2)
Note that $\|(I-F_{\xi})A(I-F_{\xi})\|\leq\|A\|=1$. Considering the restriction of $(I-F_{\xi})A(I-F_{\xi})$ on $\operatorname{ran}(I-F_{\xi})$, we have
$\sigma((I-F_{\xi})A(I-F_{\xi}))\subseteq[0,1].$ (4.3)
By (4.2) and (4.3), we have that $1+\frac{\varepsilon}{4}\notin[0,1]$. It follows that $1+\frac{\varepsilon}{4}$ is an isolated eigenvalue of $A+K$ with multiplicity $1$. This completes the proof. ∎
4.2. Perturbation of single generators of properly infinite von Neumann algebras
Recall that, given an operator $T$ and $\varepsilon>0$, by Lemma 4.1.2 there exists a self-adjoint trace-class operator $K$ with $\|K\|_{1}<\varepsilon$ such that $\operatorname{Re}T+K$ and $\operatorname{Im}T$ can be expressed as in (1.3). In this subsection, we consider a class of operators $T$ in $\mathcal{B}(\mathcal{H})$ with its real part $\operatorname{Re}T$ and imaginary part $\operatorname{Im}T$ in the form
$\operatorname{Re}T:=\begin{pmatrix}\alpha&0\\ 0&A_{\infty}\end{pmatrix}\begin{array}[]{l}\operatorname{ran}E\\ \operatorname{ran}(I-E)\end{array}\quad\text{ and }\quad\operatorname{Im}T:=\begin{pmatrix}\beta&\xi^{*}\\ \xi&B_{\infty}\end{pmatrix}\begin{array}[]{l}\operatorname{ran}E\\ \operatorname{ran}(I-E)\end{array},$ (4.4)
where
  1. $(1)$
    $\alpha$ is an isolated eigenvalue of $\operatorname{Re}T$ with multiplicity $1$, and $\beta\in\mathbb{R}$;
  2. $(2)$
    $E$ is the rank-one spectral projection for $\operatorname{Re}T$ corresponding to $\{\alpha\}$;
  3. $(3)$
    $\xi\in\operatorname{ran}(I-E)$ is in a column form and $\xi^{*}$ is the conjugate vector of $\xi$ in a row form;
  4. $(4)$
    $A_{\infty}$ and $B_{\infty}$ are self-adjoint operators on $\operatorname{ran}(I-E)$ such that $W^{*}(A_{\infty}+iB_{\infty})$ is properly infinite.
Recall that a von Neumann algebra $\mathcal{M}$ is said to be properly infinite if the identity operator $I$ is properly infinite in $\mathcal{M}$, which is equivalent to say that each central projection of $\mathcal{M}$ is either zero or infinite. The reader is referred to [14, Definition 6.3.1] for more details. By applying [14, Lemma 6.3.3], we can obtain that each properly infinite von Neumann algebra $\mathcal{M}$ can be written as a tensor product von Neumann algebra
=()¯$\mathcal{M}=\mathcal{B}(\mathcal{H}_{\infty})\overline{\otimes}\mathcal{N},$ (4.5)
where $\mathcal{H}_{\infty}$ is an infinite dimensional separable Hilbert space and $\mathcal{N}$ is a von Neumann algebra. To be self-contained, we sketch a proof for the representation (4.5) by the definitions of a properly infinite von Neumann algebra and Murray-von Neumann equivalence of projections.
For two projections $P$ and $Q$ in $\mathcal{M}$, if there exists a partial isometry $V$ in $\mathcal{M}$ such that $V^{*}V=P$ and $VV^{*}=Q$ then $P$ and $Q$ are said to be Murray-von Neumann equivalent and we denote by $P\sim Q$ this equivalence relation.
Lemma 4.2.1.
Let $\mathcal{M}$ be a properly infinite von Neumann algebra. Then there is a system of matrix units $\{E_{jk}\}_{j,k=1}^{\infty}$ in $\mathcal{M}$ such that $\sum_{j=1}^{\infty}E_{jj}=I$.
Proof.
By [14, Lemma 6.3.3], there are projections $P_{1},Q_{1}$ such that $I=P_{1}+Q_{1}$ and $P_{1}\sim Q_{1}\sim I$. Similarly, there are projections $P_{2},Q_{2}$ such that $Q_{1}=P_{2}+Q_{2}$ and $P_{2}\sim Q_{2}\sim I$. Inductively, we can define $P_{n},Q_{n}$ such that $Q_{n-1}=P_{n}+Q_{n}$ and $P_{n}\sim Q_{n}\sim I$. Let
$E_{1}=P_{1}+(I-P_{2}-P_{3}-\cdots),\quad E_{2}=P_{2},\quad E_{3}=P_{3},\quad\ldots.$
Then $I=\sum_{j=1}^{\infty}E_{j}$ and $E_{j}\sim I$ for each $j$. Let $E_{j1}$ be a partial isometry such that $E_{j1}^{*}E_{j1}=E_{1}$ and $E_{j1}E_{j1}^{*}=E_{j}$. We define $E_{ij}:=E_{i1}E_{1j}$. Then $\{E_{jk}\}_{j,k=1}^{\infty}$ is a system of matrix units in $\mathcal{M}$ such that $\sum_{j=1}^{\infty}E_{jj}=I$. ∎
Remark 4.2.2.
Let $\{T_{j}\}^{N}_{j=1}$ be an at most countable set of operators such that each $W^{*}(T_{j})$ is a finite von Neumann algebra acting on $\mathcal{H}_{j}$. Clearly, $\bigoplus^{N}_{j=1}W^{*}(T_{j})$ is a finite von Neumann algebra by applying Lemma $6.3.6$ of [14]. Note that $W^{*}(\bigoplus^{N}_{j=1}T_{j})$ is a von Neumann subalgebra of $\bigoplus^{N}_{j=1}W^{*}(T_{j})$. Employing Proposition $6.3.2$ of [14], we obtain that $W^{*}(\bigoplus^{N}_{j=1}T_{j})$ is a finite von Neumann algebra acting on $\bigoplus^{N}_{j=1}\mathcal{H}_{j}$.
In view of Remark 4.2.2, we prove an analogous result for operators $\{T_{j}\}^{N}_{j=1}$, where each $W^{*}(T_{j})$ is a properly infinite von Neumann algebra.
Lemma 4.2.3.
Let $\{T_{j}\}^{N}_{j=1}$ be an at most countable set of operators such that each $W^{*}(T_{j})$ is a properly infinite von Neumann algebra acting on $\mathcal{H}_{j}$. Then $W^{*}(\bigoplus_{j=1}^{N}T_{j})$ is a properly infinite von Neumann algebra acting on $\bigoplus^{N}_{j=1}\mathcal{H}_{j}$.
Proof.
Without loss of generality, each $\mathcal{H}_{j}$ is also viewed as a subspace of $\mathcal{H}=\bigoplus_{j=1}^{N}\mathcal{H}_{j}$. Let $E_{j}^{\prime}$ be the projection from $\mathcal{H}$ onto $\mathcal{H}_{j}$ for all $j\geq 1$.
Write $T:=\bigoplus_{j=1}^{N}T_{j}$. It is clear that $\{E_{j}^{\prime}\}_{j=1}^{N}$ is an at most countable set of projections in $W^{*}(T)^{\prime}$ with sum $I$. Let $P$ be a finite central projection in $W^{*}(T)$. We prove that $P=0$ as follows. By applying Proposition 6.3.2 of [14], we obtain that $PC_{E_{j}^{\prime}}$ is a finite central projection in $W^{*}(TC_{E_{j}^{\prime}})$, where $C_{E_{j}^{\prime}}$ is the central support of $E_{j}^{\prime}$ in $W^{*}(T)^{\prime}$ for each positive integer $j$. Furthermore, by Proposition 5.5.5 of [13], $PE_{j}^{\prime}$ is a finite central projection in $W^{*}(TE_{j}^{\prime})=W^{*}(T_{j})$. Since each $W^{*}(T_{j})$ is properly infinite, we see that $PE_{j}^{\prime}=0$ for every $j\geq 1$. It follows that $P=0$. This completes the proof. ∎
A vector $\xi$ in $\mathcal{H}$ is generating for a von Neumann algebra $\mathcal{M}$ if the set $\mathcal{M}\xi$ is dense in $\mathcal{H}$. To perturb an operator $T$ as in (4.4) to be irreducible in the trace-class norm, we need the following characterization of properly infinite von Neumann algebras with respect to generating vectors.
Lemma 4.2.4.
Let $\mathcal{H}$ be a complex separable infinite-dimensional Hilbert space and let $\mathcal{M}$ be a properly infinite von Neumann algebra acting on $\mathcal{H}$. Then the set of generating vectors of $\mathcal{M}$ is dense in $\mathcal{H}$.
Proof.
Since $\mathcal{M}$ is properly infinite, by applying Lemma 4.2.1, there is a system of matrix units $\{E_{jk}\}_{j,k=1}^{\infty}$ in $\mathcal{M}$ such that $\sum_{j=1}^{\infty}E_{jj}=I$. Let $\mathcal{N}=E_{11}\mathcal{M}E_{11}\subseteq\mathcal{B}(E_{11}\mathcal{H})$. We define a unitary operator $U\colon\mathcal{H}\to\ell^{2}\otimes E_{11}\mathcal{H}$ by
$U\xi=\sum_{j=1}^{\infty}e_{j}\otimes E_{1j}\xi,$
where $\{e_{j}\}_{j=1}^{\infty}$ is an orthonormal basis for $\ell^{2}$. We write $F_{jk}:=e_{j}\hat{\otimes}e_{k}$ for all $j,k\geq 1$. Then $\{F_{jk}\}^{\infty}_{j,k=1}$ is a system of matrix units for $\mathcal{B}(\ell^{2})$ satisfying
$UE_{jk}U^{*}=F_{jk}\otimes I_{\mathcal{N}}\quad\text{ for all }\ j,k\geq 1.$
It is routine to verify that $U\mathcal{M}U^{*}=\mathcal{B}(\ell^{2})\overline{\otimes}\mathcal{N}$. Without loss of generality, we assume that
$\mathcal{M}=\mathcal{B}(\ell^{2})\overline{\otimes}\mathcal{N},\quad\mathcal{H}=\ell^{2}\otimes\mathcal{H}_{0},\quad\mathcal{N}\subseteq\mathcal{B}(\mathcal{H}_{0}).$
Let $\xi=\sum_{j=1}^{\infty}e_{j}\otimes\xi_{j}\in\ell^{2}\otimes\mathcal{H}_{0}$ and $\varepsilon>0$. Then there is a sufficiently large integer $n$ such that $\sum_{j=n+1}^{\infty}\|\xi_{j}\|^{2}_{2}<\varepsilon^{2}/4$. Let $\{f_{j}\}_{j=1}^{\infty}$ be an orthonormal basis for $\mathcal{H}_{0}$ and construct a vector $\eta$ in the form
$\eta:=\sum_{j=1}^{n}e_{j}\otimes\xi_{j}+\sum_{k=1}^{\infty}\frac{\varepsilon}{2^{k+1}}e_{n+k}\otimes f_{k}.$
Then $\|\xi-\eta\|_{2}<\varepsilon$. Moreover, for every $j,k\geq 1$, we have
$e_{j}\otimes f_{k}=\frac{2^{k+1}}{\varepsilon}(F_{j,n+k}\otimes I_{\mathcal{N}})\eta\in\mathcal{M}\eta.$
Thus, $\eta$ is a generating vector for $\mathcal{M}$. ∎
By Lemma 4.2.4, there exist many generating vectors for a properly infinite von Neumann algebra $\mathcal{M}$ acting on a complex separable infinite-dimensional Hilbert space $\mathcal{H}$. The following proposition is a related application about generating vectors.
Proposition 4.2.5.
Let $\mathcal{M}$ be a von Neumann algebra acting on $\mathcal{H}$ with a generating vector $\xi$. Then
$W^{*}(\mathcal{M},\xi\hat{\otimes}\xi)=\mathcal{B}(\mathcal{H}).$
Proof.
Note that for every $T_{1}$ and $T_{2}$ in $\mathcal{M}$, we have
$(T_{1}\xi)\hat{\otimes}(T_{2}\xi)=T_{1}(\xi\hat{\otimes}\xi)T^{*}_{2}\in W^{*}(\mathcal{M},\xi\hat{\otimes}\xi).$
Since $\xi$ is a generating vector for $\mathcal{M}$, the set $\mathcal{M}\xi$ is dense in $\mathcal{H}$. It follows that the weak-operator closure of $\operatorname{span}\{(T_{1}\xi)\hat{\otimes}(T_{2}\xi):\text{for all }T_{1},T_{2}\in\mathcal{M}\}$ equals $\mathcal{B}(\mathcal{H})$. This completes the proof. ∎
By applying Lemma 3.1.1, we prove the following result, which plays an important role in the proof of Theorem 1.3.2.
Proposition 4.2.6.
Let $\mathcal{H}=\mathcal{H}_{1}\oplus\mathcal{H}_{2}$ be a complex infinite dimensional separable Hilbert space with $\mathcal{H}_{1}\neq\{0\}$. Suppose that $A$ and $B$ are self-adjoint operators in $\mathcal{B}(\mathcal{H})$ of the form
$A:=\begin{pmatrix}A_{1}&0\\ 0&A_{2}\end{pmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\end{array}\quad\text{ and }\quad B:=\begin{pmatrix}B_{11}&B_{12}\\ B^{*}_{12}&B_{22}\end{pmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\end{array}$ (4.6)
satisfying
  1. $(1)$
    $A_{1}$ is a diagonal operator with respect to an orthonormal basis $\{e_{j}\}^{N}_{j=1}$ of $\mathcal{H}_{1}$, where $N$ may be $\infty$,
  2. $(2)$
    the set of generating vectors for $W^{*}\left(A_{2}+iB_{22}\right)$ is dense in $\mathcal{H}_{2}$.
Then for every $\varepsilon>0$, there is a trace-class operator $K$ on $\mathcal{H}$ with $\|K\|_{1}\leq\varepsilon$ such that the operator $(A+iB)+K$ is irreducible in $\mathcal{B}(\mathcal{H})$.
Proof.
Applying Lemma 3.1.1 to the diagonal self-adjoint operator $A_{1}$ and the at most countable set $\Sigma=\sigma_{p}(A_{2})$, we obtain a self-adjoint trace-class operator $K_{A}$ in $\mathcal{B}(\mathcal{H})$ of the form
$K_{A}:=\begin{pmatrix}K_{1}&0\\ 0&0\end{pmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\end{array}$
such that
  1. $(1)$
    $\|K_{A}\|_{1}<\varepsilon/3$,
  2. $(2)$
    $\operatorname{ker}K_{1}=\{0\}$, i.e., $\operatorname{ker}K_{A}=\mathcal{H}_{2}$,
  3. $(3)$
    $\sigma_{p}(A_{1}+K_{1})\cap\sigma_{p}(A_{2})=\varnothing$,
  4. $(4)$
    each rank-one projection $e_{j}\hat{\otimes}e_{j}$ (see (2.1) for the definition) is in $W^{*}(A+K_{A})$.
Thus $P_{2}:=I-\sum_{j\geq 1}e_{j}\hat{\otimes}e_{j}$ is the projection onto $\mathcal{H}_{2}$ and $P_{2}\in W^{*}(A+K_{A})$ by Remark 3.1.2. Choose a sequence of positive numbers $\{\delta_{j}\}^{N-1}_{j=1}$ such that
$\delta_{j}\in(0,\varepsilon/2^{j+2})\quad\text{ and }\quad\delta_{j}+(e_{j}\hat{\otimes}e_{j})B_{11}(e_{j+1}\hat{\otimes}e_{j+1})\neq 0,\quad\text{ for each }1\leq j\leq N-1.$
Define $K_{2}:=\sum^{N-1}_{j=1}\delta_{j}(e_{j}\hat{\otimes}e_{j+1}+e_{j+1}\hat{\otimes}e_{j})$ and we have that $\|K_{2}\|_{1}<\varepsilon/3$. It is routine to verify that $(A_{1}+K_{1})+i(B_{11}+K_{2})$ is irreducible on $\mathcal{H}_{1}$, since $W^{*}(A_{1}+K_{1},B_{11}+K_{2})$ contains the system of matrix units $\{e_{j}\hat{\otimes}e_{k}\}^{\infty}_{j,k=1}$ generating $\mathcal{B}(\mathcal{H}_{1})$ as a von Neumann algebra.
Given a unit vector $\eta$ in $\mathcal{H}_{1}$, by the hypothesis that the set of generating vectors for $W^{*}\left(A_{2}+iB_{22}\right)$ is dense in $\mathcal{H}_{2}$, there is a vector $\xi$ in $\mathcal{H}_{2}$ with $\|\xi\|<\varepsilon/6$ such that $B^{*}_{12}\eta+\xi$ is a generating vector for $W^{*}\left(A_{2}+iB_{22}\right)$. Thus $B^{*}_{12}\eta+\xi$ is a separating vector for $W^{*}\left(A_{2}+iB_{22}\right)^{\prime}$. Define a self-adjoint trace-class operator $K_{B}$ in the form
$K_{B}:=\begin{pmatrix}K_{2}&(\xi\hat{\otimes}\eta)^{*}\\ \xi\hat{\otimes}\eta&0\end{pmatrix}.$
It is routine to verify that $\|K_{B}\|_{1}<2\varepsilon/3$ by (4.1) and the inequality $\|K_{B}\|_{1}\leq\|K_{2}\|_{1}+2\|\xi\hat{\otimes}\eta\|_{1}$. Thus we have that $(A+K_{A})+i(B+K_{B})$ is in the form
$(A+K_{A})+i(B+K_{B})=\begin{pmatrix}A_{1}+K_{1}&0\\ 0&A_{2}\end{pmatrix}+i\begin{pmatrix}B_{11}+K_{2}&B_{12}+\eta\hat{\otimes}\xi\\ B^{*}_{12}+\xi\hat{\otimes}\eta&B_{22}\end{pmatrix}.$
Let $Q$ commute with $(A+K_{A})+i(B+K_{B})$. Then we have $Q(A+K_{A})=(A+K_{A})Q$. Since $P_{2}\in W^{*}(A+K_{A})$, it follows that $QP_{2}=P_{2}Q$ and we can write $Q$ in the form
$Q=\begin{pmatrix}Q_{1}&0\\ 0&Q_{2}\end{pmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\end{array}.$
This yields that $Q_{2}$ commutes with $A_{2}+iB_{22}$. Since $Q_{1}$ commutes with the irreducible operator $(A_{1}+K_{1})+i(B_{11}+K_{2})$ on $\mathcal{H}_{1}$, we have that $Q_{1}$ is trivial on $\mathcal{H}_{1}$. Assume that $Q_{1}={0}$ and $Q={0}\oplus Q_{2}$. Since $Q(B+K_{B})=(B+K_{B})Q$, we obtain that $Q_{2}(B^{*}_{12}+\xi\hat{\otimes}\eta)=0$. Thus
$Q_{2}(B^{*}_{12}+\xi\hat{\otimes}\eta)\eta=0\Longrightarrow Q_{2}(B^{*}_{12}\eta+\xi)=0.$
Note that $(B^{*}_{12}\eta+\xi)$ is a separating vector for $W^{*}(A_{2}+iB_{22})^{\prime}$. Therefore, we have $Q_{2}={0}$ and $Q=0$. This means that $(A+iB)+(K_{A}+iK_{B})$ is irreducible in $\mathcal{B}(\mathcal{H})$. Write $K:=K_{A}+iK_{B}$. We have $\|K\|_{1}<\varepsilon$. This completes the proof. ∎
By applying Lemma 4.2.4 and Proposition 4.2.6, since $W^{*}(A_{\infty}+iB_{\infty})$ is properly infinite, we obtain that an operator $T$ as in (4.4) can be irreducible in $\mathcal{B}(\mathcal{H})$ up to an arbitrarily small trace-class perturbation.
Proposition 4.2.7.
Suppose that an operator $T=A+iB$ is in the form of (4.4), i.e., $A$ and $B$ are self-adjoint operators in the form
$A=\begin{pmatrix}\alpha&0\\ 0&A_{\infty}\end{pmatrix}\begin{array}[]{l}\operatorname{ran}E\\ \operatorname{ran}(I-E)\end{array}\quad\text{ and }\quad B=\begin{pmatrix}\beta&\xi^{*}\\ \xi&B_{\infty}\end{pmatrix}\begin{array}[]{l}\operatorname{ran}E\\ \operatorname{ran}(I-E)\end{array},$
where $\alpha$ is an isolated eigenvalue of $A$ with multiplicity $1$ and $E$ is the rank-one spectral projection for $A$ corresponding to $\{\alpha\}$. Also, assume that $W^{*}(A_{\infty}+iB_{\infty})$ is properly infinite.
Then for every $\varepsilon>0$, there exists a self-adjoint trace-class operator $K$ such that
$\|K\|_{1}<\varepsilon\quad\text{ and }\quad W^{*}(T+iK)=\mathcal{B}(\mathcal{H}).$
5. Perturbation of single generators of type $\mathrm{II}_{1}$ von Neumann algebras in trace-class norm.
In this section, we consider a class of operators $T$ in $\mathcal{B}(\mathcal{H})$ satisfying
$\operatorname{Re}T:=\begin{pmatrix}\alpha&0\\ 0&A_{2}\end{pmatrix}\begin{array}[]{l}\operatorname{ran}E\\ \operatorname{ran}(I-E)\end{array}\quad\text{ and }\quad\operatorname{Im}T:=\begin{pmatrix}\beta&\xi^{*}\\ \xi&B_{2}\end{pmatrix}\begin{array}[]{l}\operatorname{ran}E\\ \operatorname{ran}(I-E)\end{array},$ (5.1)
where
  1. $(1)$
    $\alpha$ is an isolated eigenvalue of $\operatorname{Re}T$ with multiplicity $1$ and $\beta\in\mathbb{R}$;
  2. $(2)$
    $E$ is the rank-one spectral projection for $\operatorname{Re}T$ corresponding to $\{\alpha\}$;
  3. $(3)$
    On $\operatorname{ran}(I-E)$, the von Neumann algebra $W^{*}(A_{2}+iB_{2})$ is of type $\mathrm{II}_{1}$.
In the following subsection, we discuss several cases of $T$ with respect to $W^{*}(A_{2}+iB_{2})$ satisfying different conditions.
5.1. The case for $W^{*}(A_{1}+iB_{1})$ being a type $\mathrm{II}_{1}$ von Neumann algebra $(\mathcal{M},\tau)$ acting on $L^{2}(\mathcal{M},\tau)P$.
In the following lemma, we need to apply the standard form of a finite von Neumann algebra $\mathcal{M}$ based on the GNS construction. Assume that $\mathcal{M}$ is a finite von Neumann algebra with a faithful normal tracial state $\tau$. We denote by $L^{2}(\mathcal{M},\tau)$ the underlying Hilbert space. Write $\pi_{\tau}$ as the GNS representation associated with $\tau$. Since $\tau$ is faithful, every operator $X$ in $\mathcal{M}$ can be viewed as a vector $\widehat{X}$ in $L^{2}(\mathcal{M},\tau)$ and the inner product on $\mathcal{M}$ (as a dense subset of $L^{2}(\mathcal{M},\tau)$) is defined by
$\langle\widehat{X},\widehat{Y}\rangle:=\tau(Y^{*}X),\quad\forall X,Y\in\mathcal{M}.$
In particular, we write $\hat{1}$ as the vector in $L^{2}(\mathcal{M},\tau)$ corresponding to the identity operator $I$ in $\mathcal{M}$. Recall that the representation $\pi_{\tau}$ is a normal * isomorphism satisfying
$\pi_{\tau}(T)\widehat{X}:=\widehat{TX}\quad\text{and}\quad\tau(T):=\langle\pi_{\tau}(T)\hat{1},\hat{1}\rangle,\quad\forall\,T,X\in\mathcal{M}.$
When no confusion can arise, we will omit $\pi_{\tau}$ for each operator $T$ in $\mathcal{M}$ acting on $L^{2}(\mathcal{M},\tau)$. Thus, for every $X\in\mathcal{M}$, we also write the vector $\widehat{X}$ as $X\hat{1}$.
Recall that $J$ is a conjugate-linear isometric mapping from $L^{2}(\mathcal{M},\tau)$ onto $L^{2}(\mathcal{M},\tau)$ satisfying
$J(\widehat{X}):=\widehat{X^{*}}\quad\text{ for every }X\in\mathcal{M}.$
It is clear that $J^{2}=I$. The operator $J$ is referred to as the modular conjugation operator for $(\mathcal{M},\tau)$.
With respect to the standard form of $\mathcal{M}$, there are two properties applied in this section:
  1. $(1)$
    the Hilbert space $L^{2}(\mathcal{M},\tau)$ contains the vector $\hat{1}$ which is simultaneously cyclic and separating for $\mathcal{M}$.
  2. $(2)$
    the commutant $\mathcal{M}^{\prime}$ of $\mathcal{M}$ acting on $L^{2}(\mathcal{M},\tau)$ satisfies $\mathcal{M}^{\prime}=J\mathcal{M}J$. In this sense, the Hilbert space $L^{2}(\mathcal{M},\tau)$ can be viewed as an $\mathcal{M}$$-\mathcal{M}$-bimodule, with the right action of $\mathcal{M}$ being given by
    $(Y\hat{1})\cdot X:=(JX^{*}J)(Y\hat{1})=YX\hat{1},\quad\text{ for every }X,Y\in\mathcal{M}.$ (5.2)
    The extension of the equality (5.2) implies $\xi\cdot X:=JX^{*}J\xi$ for every $X\in\mathcal{M}$ and $\xi\in L^{2}(\mathcal{M},\tau)$.
For more details, the reader is referred to [11, Section 1.2] or [14, Section 9.2].
Furthermore, for $(\mathcal{M},\tau)$ being a type $\mathrm{II}_{1}$ factor, we can view a general underlying Hilbert space $\mathcal{H}$ as an $\mathcal{M}$-$\mathbb{M}_{\infty}(\mathcal{M})$-bimodule, where $\mathbb{M}_{\infty}(\mathcal{M}):=\mathcal{M}\overline{\otimes}\mathcal{B}(\ell^{2})$ is a type $\mathrm{II}_{\infty}$ factor and the “faithful normal semifinite trace” $\mathrm{Tr}$ on $\mathbb{M}_{\infty}(\mathcal{M})$ is given by $\mathrm{Tr}\big{(}(P_{jk})_{j,k\geq 1}\big{)}=\sum^{\infty}_{j=1}\tau(P_{jj})$ for $(P_{jk})_{j,k\geq 1}$ in $\mathbb{M}_{\infty}(\mathcal{M})$. The actions of $\mathcal{M}$ and $\mathbb{M}_{\infty}(\mathcal{M})$ on $\mathcal{H}$ is defined in the following way: for $\xi=\left(\xi_{1},\xi_{2},\ldots\right)$, $X\in\mathcal{M}$, $Y=\left(Y_{jk}\right)_{j,k\geq 1}\in\mathbb{M}_{\infty}(\mathcal{M})$, and $\eta=X\xi Y$, we have $\eta_{k}=\sum_{j\geq 1}X\xi_{j}Y_{jk}$. In this terminology, we cite [11, Theorem 2.2.2] as follows.
Proposition 5.1.1 (Theorem 2.2.2 of [11]).
Let $(\mathcal{M},\tau)$ be a type $\mathrm{II}_{1}$ factor. If $\mathcal{H}$ is any separable $\mathcal{M}$-module, then there is a projection $P$ in $\mathbb{M}_{\infty}(\mathcal{M})$ such that $\mathcal{H}\cong\big{(}L^{2}(\mathcal{M},\tau)\otimes\ell^{2}\big{)}P$, and such a projection $P$ is determined uniquely up to Murray-von Neumann equivalence.
Proposition 5.1.2.
Assume that $A$ and $B$ are self-adjoint operators in $\mathcal{B}(\mathcal{H})$ in the form
$A:=\begin{pmatrix}\alpha&0\\ 0&A_{2}\end{pmatrix}\begin{array}[]{l}\operatorname{ran}E\\ \operatorname{ran}(I-E)\end{array}\quad\text{ and }\quad B:=\begin{pmatrix}\beta&\xi^{*}\\ \xi&B_{2}\end{pmatrix}\begin{array}[]{l}\operatorname{ran}E\\ \operatorname{ran}(I-E)\end{array}$ (5.3)
where
  1. $(1)$
    $\alpha$ is an isolated eigenvalue of $A$ with multiplicity $1$ and $\beta\in\mathbb{R}$;
  2. $(2)$
    $E$ is the rank-one spectral projection for $A$ corresponding to $\{\alpha\}$;
  3. $(3)$
    On $\operatorname{ran}(I-E)$, the von Neumann algebra $\mathcal{M}:=W^{*}(A_{2}+iB_{2})$ is of type $\mathrm{II}_{1}$ with a faithful normal tracial state $\tau$ and $\operatorname{ran}(I-E)=L^{2}(\mathcal{M},\tau)P$, where $P\in\mathcal{M}$ is a projection with $0<\tau(P)\leq 1$. That is to say, $\mathcal{M}$ acts on $\operatorname{ran}(I-E)$ as a left $\mathcal{M}$ module.
Then for every $\varepsilon>0$ there exists a self-adjoint trace-class operator $K$ with $\|K\|_{1}\leq\varepsilon$ such that $A+i(B+K)$ is irreducible on $\mathcal{H}$.
Proof.
It is sufficient to prove that the set of generating vectors for $\mathcal{M}$ is dense in $L^{2}(\mathcal{M},\tau)P$. Note that $\mathcal{M}P$ can be viewed as a dense subset of $L^{2}(\mathcal{M},\tau)P$. In this view, for each $\xi\in L^{2}(\mathcal{M},\tau)P$ and $\varepsilon>0$, there exists a vector $\eta\in\mathcal{M}P$ such that
$\|\xi-\eta\|_{2}=\tau\big{(}(\xi-\eta)^{*}(\xi-\eta)\big{)}^{\frac{1}{2}}<\frac{\varepsilon}{2}.$
On the other hand, consider $\mathcal{M}P$ as a subset of $\mathcal{M}$ and write the polar decomposition of $\eta$ as $\eta=\omega|\eta|$. By applying [14, Exercise 6.9.6], we assume that $\omega$ is a unitary operator in $\mathcal{M}$.
Write $\zeta:=\omega(|\eta|+\dfrac{\varepsilon}{2}{P})$. It follows that
$\|\xi-\zeta\|_{2}=\|\xi-\omega(|\eta|+\frac{\varepsilon}{2}{P})\|_{2}\leq\left\|\xi-\eta\right\|_{2}+\frac{\varepsilon}{2}\tau(P)^{\frac{1}{2}}<\varepsilon.$
Since $|\eta|+\dfrac{\varepsilon}{2}{P}$ can be viewed as an invertible operator in $P\mathcal{M}P$, we obtain that $\mathcal{M}\zeta=\mathcal{M}P$. Thus $\zeta$ is a generating vector for $\mathcal{M}$. This yields that the set of generating vectors for $\mathcal{M}$ is dense in $L^{2}(\mathcal{M},\tau)P$. Therefore, by applying Proposition 4.2.6, we complete the proof. ∎
In the following remark, we provide a useful example with respect to generating vectors and type $\mathrm{II}_{1}$ von Neumann algebras.
Remark 5.1.3.
Let $(\mathcal{M},\tau)$ be a type $\mathrm{II}_{1}$ von Neumann algebra with a faithful normal tracial state $\tau$. For a nonzero projection $P$ in $(\mathcal{M},\tau)$, the vector $\widehat{P}$ is a generating vector in $L^{2}(\mathcal{M},\tau)P$ for $\mathcal{M}$. As an application of Proposition 4.2.5, we have
$W^{*}(\mathcal{M},\widehat{P}\hat{\otimes}\widehat{P})=\mathcal{B}\big{(}L^{2}(\mathcal{M},\tau)P\big{)}.$
5.2. The case for $W^{*}(A_{2}+iB_{2})$ being a type $\mathrm{II}_{1}$ factor acting on a Hilbert space.
It is worth mentioning that if the subspace $\operatorname{ran}(I-E)=L^{2}(\mathcal{M},\tau)P$ with a nonzero projection $P\in\mathcal{M}$ in Proposition 5.1.2 is changed with $\operatorname{ran}(I-E)=\big{(}L^{2}(\mathcal{M},\tau)\otimes\ell^{2}\big{)}P^{\prime}$ for a nonzero projection $P^{\prime}\in\mathbb{M}_{\infty}(\mathcal{M})$, then the above proof method fails to work. In this subsection, for an operator $T$ as mentioned in (5.1), we consider the case for $W^{*}(A_{2}+iB_{2})$ being a type $\mathrm{II}_{1}$ factor acting on a general Hilbert space, and our goal is to express $A_{2}+iB_{2}$ as a direct sum of at most countably many irreducible operators up to an arbitrarily small trace-class perturbation. In the following Proposition 5.2.2, we show that if $A_{2}$ contains a scalar part, then $A_{2}+iB_{2}$ can be expressed as an irreducible operator up to an arbitrarily small trace-class perturbation.
The following lemma is prepared for Proposition 5.2.2. Recall that, for an operator $X$ in a von Neumann algebra $\mathcal{M}$ on $\mathcal{H}$, denote by $R(X)$ the range projection of $X$.
Lemma 5.2.1.
Let $(\mathcal{M},\tau)$ be a type $\mathrm{II}_{1}$ factor. Suppose that $B$ is a self-adjoint operator in $\mathcal{M}$ and $P$ is a nonzero projection in $\mathcal{M}$. Then there is a system of matrix units $\{E_{jk}\}_{j,k=1}^{n}$ such that $E_{11}\leq P$ and $E_{jj}BE_{kk}=0$ for all $|k-j|>1$.
Proof.
Let $P_{1}$ be a subprojection of $P$ with $\tau(P_{1})={1}/{n}$ for some integer $n\geq 1$. Since $R((I-P_{1})BP_{1})$ is a subprojection of $I-P_{1}$ with $\tau(R((I-P_{1})BP_{1}))\leq 1/n$, there is a projection $P_{2}$ with $\tau(P_{2})={1}/{n}$ such that
${R}((I-P_{1})BP_{1})\leq P_{2}\leq I-P_{1}.$
Similarly, there is a projection $P_{3}$ with $\tau(P_{3})={1}/{n}$ such that
${R}((I-P_{1}-P_{2})BP_{2})\leq P_{3}\leq I-P_{1}-P_{2}.$
Inductively, for $2\leq j\leq n$, there is a projection $P_{j}$ with $\tau(P_{j})={1}/{n}$ such that
${R}((I-P_{1}-P_{2}-\cdots-P_{j-1})BP_{j-1})\leq P_{j}\leq I-P_{1}-P_{2}-\cdots-P_{j-1}.$
Since $P_{1},P_{2},\ldots,P_{n}$ are orthogonal projections with trace ${1}/{n}$, there is a system of matrix units $\{E_{jk}\}_{j,k=1}^{n}$ in $\mathcal{M}$ such that $E_{jj}=P_{j}$ for $1\leq j\leq n$. For $k>j+1$, we have
$P_{k}BP_{j}=P_{k}(I-P_{1}-P_{2}-\cdots-P_{j})BP_{j}=P_{k}P_{j+1}(I-P_{1}-P_{2}-\cdots-P_{j})BP_{j}=0.$
Since $B$ is self-adjoint, we have $P_{k}BP_{j}=0$ for $|j-k|>1$. This completes the proof. ∎
Proposition 5.2.2.
Let $A,B$ be self-adjoint operators in a type $\mathrm{II}_{1}$ factor $(\mathcal{M},\tau)$ acting on $\mathcal{H}$ such that $W^{*}(A)$ is not diffuse. Then for any $\varepsilon>0$, there is a trace-class operator $K\in\mathcal{B}(\mathcal{H})$ with $\|K\|_{1}<\varepsilon$ such that $(A+iB)+K$ is irreducible.
Proof.
Let $P$ be a minimal projection in $W^{*}(A)$ and $\{E_{jk}\}_{j,k=1}^{n}$ the system of matrix units given by Lemma 5.2.1 with $E_{11}\leq P$.
Note that $E_{11}A=AE_{11}$ is a scalar multiple of $E_{11}$. By applying Lemma 3.1.1, there exists a self-adjoint trace-class operator $K_{0}$ on $E_{11}\mathcal{H}$ with $\|K_{0}\|_{1}<\frac{\varepsilon}{2n}$ such that $(A+K_{0})E_{11}$ is a diagonal operator on $E_{11}\mathcal{H}$ with distinct eigenvalues and the point spectrum of $(A+K_{0})E_{11}$ on $E_{11}\mathcal{H}$ is disjoint from the point spectrum of $A(I-E_{11})$ on $(I-E_{11})\mathcal{H}$. It follows that $E_{11}\in W^{*}(A+K_{0})$.
Since $(A+K_{0})E_{11}$ is diagonal on $E_{11}\mathcal{H}$ with distinct eigenvalues, there is a self-adjoint trace-class operator $K_{1}$ on $E_{11}\mathcal{H}$ with $\|K_{1}\|_{1}<\frac{\varepsilon}{2n}$ such that $(A+K_{0})E_{11}+iE_{11}(B+K_{1})E_{11}$ is irreducible on $E_{11}\mathcal{H}$.
For $2\leq j\leq n$, since $E_{11}-{R}(E_{1,j-1}BE_{j1})\sim E_{11}-{R}(E_{1j}BE_{j-1,1})$ in $\mathcal{M}$, there is a trace-class operator $K_{j}$ on $E_{11}\mathcal{H}$ with $\|K_{j}\|_{1}<\frac{\varepsilon}{2n}$ such that $E_{1,j-1}BE_{j1}+K_{j}$ has trivial kernel on $E_{11}\mathcal{H}$. Define a trace-class operator $K$ in the form
$K:=K_{0}+i\left(K_{1}+\sum_{j=2}^{n}(E_{j-1,1}K_{j}E_{1j}+E_{j1}K_{j}^{*}E_{1,j-1})\right).$
It follows that $\|K\|_{1}<\varepsilon$.
Notice that the real part of $(A+iB)+K$ is
$A+K_{0}=(A+K_{0})E_{11}+A(I-E_{11}),$
and the imaginary part of $(A+iB)+K$ is
$\begin{split}B^{\prime}&=E_{11}(B+K_{1})E_{11}+\sum_{j=2}^{n}E_{jj}BE_{jj}\\ &\quad+\sum_{j=2}^{n}E_{j-1,1}(E_{1,j-1}BE_{j1}+K_{j})E_{1j}+\sum_{j=2}^{n}E_{j1}(E_{1j}BE_{j-1,1}+K_{j}^{*})E_{1,j-1}.\end{split}$
Let $Q$ be a projection commuting with $(A+iB)+K$. We claim that $Q=0$ or $Q=I$. Note that the equality $Q(A+K_{0})=(A+K_{0})Q$ implies that $Q$ commutes with $E_{11}\in W^{*}(A+K_{0})$. We write $Q=Q_{1}+Q_{2}$, where $Q_{1}=QE_{11}$ and $Q_{2}=Q(I-E_{11})$. Since the restriction of the operator $(A+K_{0})E_{11}+iE_{11}(B+K_{1})E_{11}$ is irreducible on $E_{11}\mathcal{H}$, we have $Q_{1}=0$ or $Q_{1}=E_{11}$.
Without loss of generality, we assume that $Q_{1}=0$. Thus, $Q_{2}=Q_{1}+Q_{2}=Q$ commutes with the imaginary part $B^{\prime}$ of $(A+iB)+K$. Combining with $E_{11}Q_{2}=0$, we obtain the equality $E_{11}B^{\prime}Q_{2}=E_{11}Q_{2}B^{\prime}=0$. It follows that
$(E_{11}BE_{21}+K_{2})E_{12}Q_{2}E_{21}=E_{11}B^{\prime}Q_{2}E_{21}=0.$
Since $E_{11}BE_{21}+K_{2}$ has trivial kernel on $E_{11}\mathcal{H}$, we have $E_{12}Q_{2}E_{21}=0$, i.e., $Q_{2}E_{22}=0$.
Similarly, since $E_{12}BE_{31}+K_{3}$ has trivial kernel on $E_{11}\mathcal{H}$, we have $Q_{2}E_{33}=0$. Inductively, we can get $Q_{2}E_{jj}=0$ for $2\leq j\leq n$. Thus, $Q_{2}=0$. Therefore, $(A+iB)+K$ is irreducible. ∎
Remark 5.2.3.
Among examples of single operators $A+iB$ of type $\mathrm{II}_{1}$ factors, for each separable type $\mathrm{II}_{1}$ factor $\mathcal{M}$ with $\mathcal{G}(\mathcal{M})=0$, by Theorem $16.4.5$ of [27], it is not hard to construct two self-adjoint operators $A$ and $B$ with $W^{*}(A+iB)=\mathcal{M}$ and $W^{*}(A)$ not diffuse. It is also worth mentioning that for each property $T$ factor $\mathcal{N}$ associated with groups $SL_{n}(\mathbb{Z})$ for $n\geq 3$, by Theorem $5$ of [8], there exists two self-adjoint operators $A$ and $B$ with $W^{*}(A+iB)=\mathcal{N}$ and $W^{*}(A)$ not diffuse. An interesting question is whether $\mathcal{L}(\mathbf{F}_{2})$ has a single generator $A+iB$ with either $W^{*}(A)$ or $W^{*}(B)$ not diffuse.
In the rest part of this subsection, comparing with Proposition 5.2.2, for an operator $T$ as in (5.1) with $A_{2}+iB_{2}$ generating a type $\mathrm{II}_{1}$ factor, we only need to discuss the case that $W^{*}(A_{2})$ is diffuse.
Recall that, given a type $\mathrm{II}_{1}$ factor $\mathcal{M}$, the normalizer $\mathcal{N}(\mathcal{A})$ for a masa $\mathcal{A}$ of $\mathcal{M}$ is defined as in the form
$\mathcal{N}(\mathcal{A}):=\left\{U\in\mathcal{M}\ |\ U\mathcal{A}U^{*}=\mathcal{A},\ U\text{ unitary}\right\}.$
Moreover, $\mathcal{A}$ is said to be Cartan (or regular) if $\mathcal{N}(\mathcal{A})^{\prime\prime}=\mathcal{M}$. For more details, the reader is referred to [27, Section 3.3]. A separably acting von Neumann algebra is said to be hyperfinite if it can be expressed as the weak-operator closure of an ascending sequence of finite dimensional subalgebras. The reader is referred to [27, Section 3.8] and [14, Chapter 12] for more details.
Lemma 5.2.4.
Let $\mathcal{M}$ be a type $\mathrm{II}_{1}$ factor acting on $\mathcal{H}$ generated by two unitary operators $U$ and $V$, where $UV=e^{2\pi i\theta}VU$ and $\theta$ is an irrational number.
Then there is a sequence $\{n_{k}\}_{k=1}^{\infty}$ of increasing positive integers such that $U^{*n_{k}}TU^{n_{k}}$ is weak-operator convergent to $T$ for every $T\in\mathcal{M}$.
Proof.
Since $\theta$ is irrational, there is a sequence $\{n_{k}\}_{k=1}^{\infty}$ of increasing positive integers such that
$|e^{2\pi i\theta mn_{k}}-1|<\frac{1}{2^{k}}\quad\text{for}\leavevmode\nobreak\ -k\leq m\leq k.$
We will show that $\{n_{k}\}_{k=1}^{\infty}$ has the desired property.
Let $\pi$ be the GNS representation of $\mathcal{M}$ on $L^{2}(\mathcal{M},\tau)$ with respect to the tracial state $\tau$ on $\mathcal{M}$. Since $\tau$ is faithful and normal, it is a *-isomorphism of $\mathcal{M}$ onto $\pi(\mathcal{M})$. Note that the sequence $\{U^{*n_{k}}TU^{n_{k}}\}_{k\geq 1}$ is bounded, by [14, Corollary 7.1.16], we only need to prove that $\pi(U^{*n_{k}}TU^{n_{k}})$ is weak-operator convergent to $\pi(T)$ for every $T\in\mathcal{M}$.
Without loss of generality, we assume that $\mathcal{H}=L^{2}(\mathcal{M},\tau)$. In this case, it clear that the set $\{U^{m}V^{n}\hat{1}\colon m,n\in\mathbb{Z}\}$ forms an orthonormal basis for $\mathcal{H}$. For any $m,n,a,b\in\mathbb{Z}$, a direct computation shows that
$\langle(U^{*n_{k}}TU^{n_{k}}-T)U^{m}V^{n}\hat{1},U^{a}V^{b}\hat{1}\rangle=(e^{2\pi i\theta(n-b)n_{k}}-1)\langle TU^{m}V^{n}\hat{1},U^{a}V^{b}\hat{1}\rangle.$
It follows that
$\lim_{k\to\infty}\langle(U^{*n_{k}}TU^{n_{k}}-T)U^{m}V^{n}\hat{1},U^{a}V^{b}\hat{1}\rangle=0.$
Since $\{U^{*n_{k}}TU^{n_{k}}-T\}_{k\geq 1}$ is a bounded sequence, it is weak-operator convergent to 00. This completes the proof. ∎
Lemma 5.2.5.
Let $\mathcal{M}$ be a type $\mathrm{II}_{1}$ factor acting on $\mathcal{H}$. If $U$ is a Haar unitary in $\mathcal{M}$ and $K$ is a compact operator in $\mathcal{B}(\mathcal{H})$, then $U^{*n}KU^{n}$ is strong-operator convergent to 00, as $|n|\to\infty$.
Proof.
Since $K$ is compact, we only need to show that $U^{n}$ is weak-operator convergent to 00. Similar to Lemma 5.2.4, we may assume $\mathcal{H}=L^{2}(\mathcal{M},\tau)$. For any $X,Y\in\mathcal{M}$, we have
$\langle U^{n}X\hat{1},Y\hat{1}\rangle=\tau(U^{n}XY^{*})=\langle U^{n}\hat{1},YX^{*}\hat{1}\rangle.$
Since $\{U^{n}\hat{1}\colon n\in\mathbb{Z}\}$ is an orthonormal set, by Bessel’s inequality, we have
$\lim_{n\to\infty}\langle U^{n}X\hat{1},Y\hat{1}\rangle=0.$
Since $\{X\hat{1}\colon X\in\mathcal{M}\}$ is dense in $L^{2}(\mathcal{M},\tau)$ and $\{U^{n}\}_{n\in\mathbb{Z}}$ is a bounded sequence, $\{U^{n}\}_{n\in\mathbb{Z}}$ is weak-operator convergent to 00, as $|n|\to\infty$. This completes the proof. ∎
Before the following result, we briefly recall the relation between the irrational rotation algebra $\mathcal{A}_{\theta}$ and the standard form of a hyperfinite $\mathrm{II}_{1}$ factor. Let $\theta$ be an irrational number. The irrational rotation algebra $\mathcal{A}_{\theta}$ is the universal ${\rm C}^{\ast}$-algebra generated by two unitary elements $U$ and $V$ satisfying
$UV=e^{2\pi i\theta}VU.$
Since $\mathcal{A}_{\theta}$ can also be viewed as a crossed product ${\rm C}^{\ast}$-algebra, an application of Theorem 1 of [24] implies that $\mathcal{A}_{\theta}$ is amenable.
It is well known that $\mathcal{A}_{\theta}$ is simple and there exists a unique faithful tracial state $\tau$ on $\mathcal{A}_{\theta}$ (see [3, Proposition VI.1.3, Theorem VI.1.4]). The reader is referred to Chapter VI of [3] for more details and results related to $\mathcal{A}_{\theta}$. By the GNS construction, the unique faithful tracial state $\tau$ induces a $\ast$-representation $\pi$ of $\mathcal{A}_{\theta}$ on $L^{2}(\mathcal{A}_{\theta},\tau)$. For a quick check that $\pi(\mathcal{A}_{\theta})^{\prime\prime}$ is a hyperfinite type $\mathrm{II}_{1}$ factor acting on $L^{2}(\mathcal{A}_{\theta},\tau)$, the reader is referred to [28, Lemma 2.11].
Proposition 5.2.6.
Let $\mathcal{M}$ be a hyperfinite type $\mathrm{II}_{1}$ factor acting on $\mathcal{H}$ generated by two self-adjoint operators $A$ and $B$.
If $\mathcal{M}$ has a generating vector and $W^{*}(A)$ is a Cartan subalgebra of $\mathcal{M}$, then for any $\varepsilon>0$, there is a self-adjoint rank-one operator with $\|K\|_{1}<\varepsilon$ such that $A+i(B+K)$ is irreducible.
Proof.
By the arguments about the hyperfinite $\mathrm{II}_{1}$ factor preceding this proposition, there are unitary operators $U$ and $V$ in $\mathcal{M}$ with $UV=e^{2\pi i\theta}VU$ for some irrational number $\theta$ such that
$W^{*}(U,V)=\mathcal{M},\quad W^{*}(U)=W^{*}(A).$
Let $K=\frac{\varepsilon}{2}\xi\hat{\otimes}\xi$, where $\xi$ is a generating unit vector of $\mathcal{M}$. Then $\|K\|_{1}<\varepsilon$. By Lemma 5.2.4 and Lemma 5.2.5, $B$ is a weak-operator limit of $U^{*n}(B+K)U^{n}$. It follows that
$B\in W^{*}(U,B+K)=W^{*}(A+i(B+K)).$
Hence $K\in W^{*}(A+i(B+K))$. Since $\xi$ is a generating unit vector of $\mathcal{M}$, by Proposition 4.2.5, we obtain that $A+i(B+K)$ is irreducible on $\mathcal{H}$. This completes the proof. ∎
Proposition 5.2.7.
Assume that $A$ and $B$ are self-adjoint operators generating a hyperfinite type $\mathrm{II}_{1}$ factor $(\mathcal{M},\tau)$ with a faithful normal tracial state $\tau$ such that $W^{*}(A)$ is a Cartan subalgebra of $(\mathcal{M},\tau)$.
Then for every $\varepsilon>0$, there exists a self-adjoint trace-class operator $K$ with $\|K\|_{1}\leq\varepsilon$ such that $A+i(B+K)$ is a direct sum of at most countably many irreducible operators.
Proof.
Let $\mathcal{H}$ be the underlying complex separable Hilbert space of $\mathcal{M}$. Note that $\mathcal{M}$ acts on $\mathcal{H}$ and every $\ast$-representation is a direct sum of cyclic representations. Thus, there is an orthogonal family $\{E^{\prime}_{j}\}$ of cyclic projections in $\mathcal{M}^{\prime}$ with sum $I$, where by a cyclic projection $E^{\prime}_{j}$ we mean $E^{\prime}_{j}\mathcal{H}=\overline{\mathcal{M}\xi_{j}}$ for some unit vector $\xi_{j}$ in $\mathcal{H}$. Since $\mathcal{H}$ is separable, we obtain at most countably many $\{\xi_{j}\}^{N}_{j=1}$.
Note that, by applying [13, Proposition 5.5.5], $W^{*}(AE^{\prime}_{j})$ is a Cartan subalgebra of the hyperfinite type $\mathrm{II}_{1}$ factor $\mathcal{M}E^{\prime}_{j}$. According to Proposition 5.2.6, there exists a self-adjoint rank-one operator $K_{j}$ with $\|K_{j}\|_{1}<\frac{\varepsilon}{2^{j}}$ such that $AE^{\prime}_{j}+i(BE^{\prime}_{j}+K_{j})$ is irreducible on $E^{\prime}_{j}\mathcal{H}$. Define $K:=\bigoplus^{N}_{j=1}K_{j}$. Hence, we have $\|K\|_{1}<\varepsilon$. Therefore
$A+i(B+K)=\bigoplus^{N}_{j=1}\left(AE^{\prime}_{j}+i(BE^{\prime}_{j}+K_{j})\right)$
is a direct sum of at most countably many irreducible operators. This completes the proof. ∎
In Proposition 5.2.7, a collection of hyperfinite type $\mathrm{II}_{1}$ factors can be viewed as a positive support of 1.3.1. It is natural to seek for more examples to support 1.3.1. In the following example, an operator generating a non-hyperfinite type $\mathrm{II}_{1}$ factor is presented.
In the following, by $\mathcal{L}(\mathbf{F}_{r})$ we denote the interpolated free group factor for $1<r\leq\infty$ (see [29, Postscript]). By applying [5, Corollary 5.7], $\mathcal{L}(\mathbf{F}_{r})$ can be generated by $\lceil r\rceil$ self-adjoint operators for $1<r\leq\infty$. Recall that for any real number $x$, let $\lfloor x\rfloor$ and $\lceil x\rceil$ be the floor function and the ceiling function of $x$, respectively.
Example 5.2.8.
Let $\mathcal{L}(\mathbf{F}_{p})$ act on a separable Hilbert space $\mathcal{H}$, where $21<p<$ 2. Then there are self-adjoint operators $A$ and $B$ generating $\mathcal{L}(F_{p})$ such that for any $\varepsilon>0$, there exists a self-adjoint trace-class operator $K$ with $\|K\|_{1}\leq\varepsilon$ such that $A+i(B+K)$ is a direct sum of irreducible operators.
Proof.
Let $n\geq 3$ be an integer such that $\lceil 1+n^{2}(p-1)\rceil\leq n^{2}$. Then $\mathcal{L}(\mathbf{F}_{1+n^{2}(p-1)})$ can be generated by $n^{2}$ positive operators
$A_{j}\ (1\leq j\leq n);\ B_{jk}\ (1\leq j\leq k\leq n,(j,k)\neq(1,1));\ C_{jk}\ (3\leq j+2\leq k\leq n).$
Furthermore, we can assume that $\sigma(A_{j})\cap\sigma(A_{k})=\varnothing$ for $j\neq k$, and each $B_{j,j+1}$ is invertible for $11\leq j\leq n-$ 1. For $1\leq j\leq k\leq n$, define
$X_{jk}:=\begin{cases}0,&\text{for }(j,k)=(1,1),\\ B_{jk}+iC_{jk},&\text{for }3\leq j+2\leq k\leq n,\\ B_{jk},&\mbox{otherwise}.\end{cases}$
Let $X_{jk}=X_{kj}^{*}$ for $1\leq k<j\leq n$. We define
$A=\mathrm{diag}(A_{1},A_{2},\ldots,A_{n}),\quad B=(X_{jk})_{1\leq j,k\leq n}.$
Then $A+iB$ is a generator of $\mathcal{L}(\mathbf{F}_{p})={\mathbb{M}}_{n}(\mathbb{C})\otimes\mathcal{L}(\mathbf{F}_{1+n^{2}(p-1)})$ (see [30, Theorem 5.1]). By considering the commutant of $\mathcal{L}(\mathbf{F}_{p})$, we may assume that $\mathcal{L}(\mathbf{F}_{p})$ acts on $\mathbb{C}^{n}\otimes\mathcal{H}_{0}$, where $\mathcal{L}(\mathbf{F}_{1+n^{2}(p-1)})$ acts on $\mathcal{H}_{0}$ and has a generating unit vector $\xi$.
Write each operator of $\mathcal{L}(\mathbf{F}_{p})$ as an $n\times n$ matrix with entries in $\mathcal{L}(\mathbf{F}_{1+n^{2}(p-1)})$, and define a self-adjoint trace-class operator $K$ to be the operator-valued matrix with the $(1,1)$-entry being $\varepsilon\xi\hat{\otimes}\xi$ on $\mathcal{H}_{0}$ and 00 at all other entries. That is to say, we can write $K$ in the form
$K:=\begin{pmatrix}\varepsilon\xi\hat{\otimes}\xi&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\end{pmatrix}.$
Note that $(\xi,0,\ldots,0)$ is the generating vector for $\mathcal{L}(\mathbf{F}_{p})$. A routine calculation ensures that $A+i(B+K)$ is irreducible on $\mathbb{C}^{n}\otimes\mathcal{H}_{0}$. ∎
With the method adopted in Example 5.2.8, we can construct a family of examples to support 1.3.1.
6. Proof of the Main Theorem
With these technical tools developed in the preceding sections, we are ready to prove Theorem 1.3.2 (Main Theorem) in the following.
Proof of Theorem 1.3.2.
Let $T$ be a reducible operator in $\mathcal{B}(\mathcal{H})$ and $\varepsilon>0$.
Step 1. By Lemma 4.1.2 and the type decomposition for von Neumann algebras, there is a self-adjoint trace-class operator $K_{0}$ with $\|K_{0}\|_{1}<\varepsilon/4$ such that $A:=\operatorname{Re}(T)+K_{0}$ and $B:=\operatorname{Im}(T)$ are of the form
$A:=\begin{pmatrix}\alpha&0&0&0\\ 0&A_{\infty}&0&0\\ 0&0&A_{{1}}&0\\ 0&0&0&A_{{2}}\\ \end{pmatrix}\begin{array}[]{l}\operatorname{ran}E\\ \mathcal{H}_{\infty}\\ \mathcal{H}_{1}\\ \mathcal{H}_{2}\\ \end{array},\quad B:=\begin{pmatrix}\beta&\xi^{*}_{0}&\xi^{*}_{1}&\xi^{*}_{2}\\ \xi_{0}&B_{\infty}&0&0\\ \xi_{1}&0&B_{{1}}&0\\ \xi_{2}&0&0&B_{{2}}\\ \end{pmatrix}\begin{array}[]{l}\operatorname{ran}E\\ \mathcal{H}_{\infty}\\ \mathcal{H}_{1}\\ \mathcal{H}_{2}\\ \end{array},$ (6.1)
where
  1. $(1)$
    $\alpha$ is an isolated eigenvalue of $A$ with multiplicity 1, and $E$ is the spectral projection of $A$ corresponding to $\{\alpha\}$;
  2. $(2)$
    $W^{*}(A_{\infty}+iB_{\infty})$ is a properly infinite von Neumann algebra in $\mathcal{B}(\mathcal{H}_{\infty})$;
  3. $(3)$
    $W^{*}(A_{1}+iB_{1})$ is a finite type I von Neumann algebra in $\mathcal{B}(\mathcal{H}_{1})$;
  4. $(4)$
    $W^{*}(A_{2}+iB_{2})$ is a type $\mathrm{II}_{1}$ von Neumann algebra in $\mathcal{B}(\mathcal{H}_{2})$.
For notational convenience, up to unitary equivalence, we rewrite matrix forms as in (1.3) as the forms in (6.1).
Step 2. If $\mathcal{H}_{1}=\{0\}$, then write $K_{1}=0$. Otherwise, by applying Proposition 3.2.7, there is a trace-class operator $K_{1}$ in $\mathcal{B}(\mathcal{H}_{1})$ with $\|K_{1}\|_{1}<\varepsilon/4$ such that $(A_{1}+iB_{1})+K_{1}$ is a direct sum of at most countably many irreducible operators. Moreover, we can require that $\alpha\notin\sigma(A_{1}+\operatorname{Re}K_{1})$.
Step 3. If $\mathcal{H}_{2}=\{0\}$, then we write $K_{2}=0$. Otherwise, by 1.3.1, there is a trace-class operator $K_{2}$ in $\mathcal{B}(\mathcal{H}_{2})$ with $\|K_{2}\|_{1}<\varepsilon/4$ such that $(A_{2}+iB_{2})+K_{2}$ is a direct sum of at most countably many irreducible operators, and $\alpha\notin\sigma(A_{2}+\operatorname{Re}K_{2})$.
If $\mathcal{H}_{1}=\{0\}$ and $\mathcal{H}_{2}=\{0\}$, then the proof is finished by Proposition 4.2.7. For the remaining part of the proof, we assume that either $\mathcal{H}_{1}\neq\{0\}$ or $\mathcal{H}_{2}\neq\{0\}$.
Step 4. With Step 2 and Step 3, we obtain that the operator of the form
$\begin{pmatrix}(A_{1}+iB_{1})+K_{1}&0\\ 0&(A_{2}+iB_{2})+K_{2}\\ \end{pmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\\ \end{array}$
is a direct sum $(\bigoplus_{j\geq 1}X_{j})\oplus(\bigoplus_{k\geq 1}Y_{k})$ of at most countably many irreducible operators, where each $\operatorname{ran}X_{j}$ is infinite dimensional and each $\operatorname{ran}Y_{k}$ is finite dimensional.
Note that the real part of $\bigoplus_{k\geq 1}Y_{k}$ is diagonal up to a unitary equivalence, and by Lemma 4.2.3, $(A_{\infty}+iB_{\infty})\oplus(\bigoplus_{j\geq 1}X_{j})$ generates a properly infinite von Neumann algebra.
If the part $(A_{\infty}+iB_{\infty})\oplus(\bigoplus_{j\geq 1}X_{j})$ vanishes, then by Lemma 2.1.2, there is a trace-class operator $K_{3}$ with $\|K_{3}\|_{1}<\varepsilon/4$ such that $T+K$ is irreducible on $\mathcal{H}$, where $K=K_{0}+K_{1}+K_{2}+K_{3}$ and $\|K\|_{1}<\varepsilon$.
Otherwise, $(A_{\infty}+iB_{\infty})\oplus(\bigoplus_{j\geq 1}X_{j})$ generates a properly infinite von Neumann algebra and $\operatorname{Re}(\bigoplus_{k\geq 1}Y_{k})\oplus\alpha E$ is diagonal. Note that $E\neq 0$. By Proposition 4.2.6, there is a trace-class operator $K_{4}$ with $\|K_{4}\|_{1}<\varepsilon/4$ such that $T+K$ is irreducible on $\mathcal{H}$, where $K=K_{0}+K_{1}+K_{2}+K_{4}$ and $\|K\|_{1}<\varepsilon$. This completes the proof. ∎